### v2

Abstract
Our interest lies in developing some efficient methods for minimizing the sum of two geodesically convex functions on Hadamard manifolds, with the aim of improving the convergence of the Douglas-Rachford algorithm in Hadamard manifolds. Specifically, we propose two types of algorithms: inertial and non-inertial algorithms. The convergence analysis of both algorithms is provided under suitable assumptions on algorithmic parameters and the geodesic convexity of the objective functions. This convergence analysis is based on fixed-point theory for nonexpansive operators. We also study the convergence rates of these two methods.
Additionally, we introduce parallel Douglas-Rachford type algorithms for minimizing functionals containing multiple summands with applications to the generalized Heron problem on Hadamard manifolds. To demonstrate the effectiveness of the proposed algorithms, we present some numerical experiments for the generalized Heron problems.
keywords:
Inertial methods, Douglas-Rachford methods, Fixed point methods, Minimization problems, Generalized Heron problems
1 Introduction
The Douglas-Rachford algorithm is among the most celebrated splitting algorithms in convex analysis and operator theory [Combettes2008Inv, Bot2015, Dixit2022, Bergmann2016, ShikherOpt2024, Hieu2018COA, Barshad2023], originally introduced by Douglas and Rachford [Douglas1956] for the numerical solution of partial differential equations. It was later generalized by Lions and Mercier [Lions1967] and independently by Passty [Passty1979] to solve the more general problem of finding a zero of the sum of two maximally monotone operators in Hilbert spaces. Eckstein and Bertsekas [Eckstein1992] provided further insights into its convergence under relaxation schemes and inexact computations.
Beyond operator inclusions, the algorithm also finds wide application in convex optimization, particularly for problems involving the sum of two proper, lower semicontinuous, and convex functions; see [Combettes2008Inv]. Given a Hilbert space $X$ and two proper, convex, lower semicontinuous functions $\phi,\psi:X\to(-\infty,+\infty]$, the Douglas–Rachford algorithm aims to solve the following convex minimization problem:
$\min_{x\in X}\phi(x)+\psi(x).$ (1)
The strength of the Douglas-Rachford algorithm lies in its splitting structure, which allows one to compute the proximal mappings of $\phi$ and $\psi$ separately. This structure makes the algorithm especially useful in large-scale and structured optimization problems [Combettes2008Inv].
Recently, these algorithms were successfully applied in image processing mainly for two reasons: the functional to minimize allows simple proximal mappings within the method and it turned out that the algorithms are highly parallelizable; see [Combettes2008Inv]. Therefore, Douglas-Rachford methods became some of the most popular variational methods for image processing. Continued interest in Douglas-Rachford iteration methods is also due to its excellent performance on various nonconvex problems; see [Bauschke2002, Borwein2011].
Motivated by applications where data naturally reside on nonlinear spaces such as manifolds, common in computer vision, machine learning, and information geometry, there has been growing interest in extending convex optimization methods to manifold settings. Hadamard manifolds, in particular, offer a natural framework for such generalizations due to their geometric properties, including unique geodesics between points [BacakNAA2014]. In the past two decades, several results from linear spaces have been extended to manifold settings because of their applications in many areas of science, engineering, and management (see [Martin2, Martin3, Ferreira2002, Fereira2005]).
[…]
$\min_{x\in\mathcal{M}}\phi(x)+\psi(x).$ (2)
To compute the solution to problem (2), Bergmann et al. [Bergmann2016] proposed the following Douglas-Rachford algorithm on Hadamard manifold:
$x_{n+1}=\gamma(x_{n},R_{\lambda\phi}R_{\lambda\psi}(x_{n});\alpha_{n}),$ (3)
where $\lambda>0$, $\{\alpha_{n}\}$ is a sequence in $[0,1]$ with $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$ and $\gamma(x,y;\cdot)$ is the unique geodesic joining $x$ to $y$ in $\mathcal{M}$. The operators $R_{\lambda\phi}$ and $R_{\lambda\psi}$ denote the reflection of $\phi$ and $\psi$, respectively. Geometrically, $R_{\lambda\phi}(x)$ is obtained by reflecting the point $x$ with respect to its proximal point $\operatorname*{Prox}_{\lambda\phi}(x)$ along the unique geodesic joining them. Bergmann et al. [Bergmann2016] showed that if $R_{\lambda\phi}R_{\lambda\psi}$ is nonexpansive, then the sequence generated by method (3) converges to an element $v$ of $\operatorname*{Fix}(R_{\lambda\phi}R_{\lambda\psi})$ such that $u=\operatorname*{Prox}_{\lambda\psi}(v)$ is solution of problem (2).
It is worth noting that by setting $T={R}_{\lambda\phi}{R}_{\lambda\psi}$, the Douglas-Rachford algorithm (3) can be seen as a Riemannian Krasnosel’skiĭ–Mann iteration of the form
$x_{n+1}=\gamma(x_{n},Tx_{n};\alpha_{n}),$
which is the natural extension of the classical Krasnosel’skiĭ–Mann scheme from Hilbert spaces to Hadamard manifolds; see [Bergmann2016]. This observation connects optimization and fixed point theory and hence problem (2) can now be viewed as a fixed point problem.
This naturally leads to the framework of fixed point theory. Various problems across diverse domains, such as image reconstruction [Byrne2004], signal processing [Byrne2004], variational inequality problems [Mercier1980], and convex feasibility problems [Bauschke1996] can be formulated as fixed point problems. Let $C$ be a nonempty, closed and convex subset of a real Hilbert space $X$ and $T:C\to C$ be a nonlinear operator. Then, the fixed point problem can be stated as:
 $\text{ find }x\in C\text{ such that }Tx=x.$ (4)
To solve such problems, numerous iterative techniques have been developed. One of the most widely used is the Mann iterative method, introduced by Mann [Mann1953] in 1953. The Mann iteration method has its limitations, as it may not effectively approximate the fixed points of pseudocontractive mappings (see [Chidume2001]). To tackle this challenge, Ishikawa [Ishikawa1974] proposed an iterative approach, extensively investigated by multiple authors across different spaces to compute the fixed point of pseudocontractive and nonexpansive mappings (see [Takahashi2008, Acedo2007, Dotson1970]).
Building on this, Agarwal et al. [Agarwal2007] introduced the S-iteration method as follows:
$\begin{cases}y_{n}=(1-\beta_{n})x_{n}+\beta_{n}Tx_{n},\\ x_{n+1}=(1-\alpha_{n})Tx_{n}+\alpha_{n}Ty_{n}\text{ for all }n\in\mathbb{N},\end{cases}$ (5)
where $\{\alpha_{n}\}$ and $\{\beta_{n}\}$ are sequences in $(0,1)$ with $\sum_{n=1}^{\infty}\alpha_{n}\beta_{n}(1-\beta_{n})=\infty$. In 2011, Sahu [Sahu2011] proposed a modified version of the S-iteration method (5), called the normal S-iteration method, also known as the hybrid Picard-Mann iteration method, which is defined as follows:
$x_{n+1}=T((1-\alpha_{n})x_{n}+\alpha_{n}Tx_{n})\text{ for all }n\in\mathbb{N},$ (6)
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$. Continuing in this direction, Sahu [Sahusoft2020] observed that applying the operator multiple times in each iteration can significantly enhance the convergence. This led to the $p$-accelerated normal S-iteration method [Sahusoft2020], defined as follows:
$x_{n+1}=T^{p}((1-\alpha_{n})x_{n}+\alpha_{n}Tx_{n})\text{ for all }n\in\mathbb{N},$ (7)
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ and $p$ is the power of acceleration. For $p=1$, the $p$-accelerated normal S-iteration method reduces to the normal S-iteration method defined by (6). It is observed in [Sahusoft2020] that the $p$-accelerated normal S-iteration method has better convergence performance than the Mann iterative method and normal S-iterative method for the fixed point of nonexpansive operators with convex domain in infinite dimensional Hilbert space. In fact in most cases $p$-accelerated normal S-iteration method works better than the inertial Mann iteration method due to the parameter $p$.
Inertial-type extrapolation techniques play a critical role in accelerating iterative methods. The foundation of such methods can be traced back to Polyak [Polyak1964], who introduced the heavy ball method. Alvarez and Attouch [Alvarez2001] extended the concept to improve the performance of the proximal point algorithm, proposing the inertial proximal point algorithm for solving inclusion problems.
In the existing literature, the utilization of inertial extrapolation-based methods has been thoroughly explored and implemented in various contexts (see [Ochs2014, Bot2016, Chen2014, Samir2023, Attouch2019] and references therein). Recently, researchers have developed several iterative algorithms by incorporating inertial extrapolation techniques, such as the fast shrinkage thresholding algorithm [Beck2009(1)], inertial forward-backward algorithm [Lorenz2015] and inertial Douglas-Rachford splitting algorithm [Bot2015].
Building upon the development of inertial and accelerated iterative methods, Mainge [Mainge2008] proposed a novel approach in 2008 by combining inertial extrapolation with the Mann method, aiming to compute the fixed point of nonexpansive mappings on a real Hilbert space. The inertial Mann algorithm is defined as follows:
$\begin{cases}y_{n}=x_{n}+\theta_{n}(x_{n}-x_{n-1}),\\ x_{n+1}=(1-\alpha_{n})y_{n}+\alpha_{n}Ty_{n}\text{ for all }n\in\mathbb{N},\end{cases}$ (8)
where $\{\theta_{n}\}$ and $\{\alpha_{n}\}$ are sequences in $[0,1)$. Mainge [Mainge2008] studied the weak convergence of inertial Mann algorithm (8) under the conditions that $\theta_{n}\in[0,\theta]$ for all $n\in\mathbb{N}$, where $\theta\in[0,1)$, $\sum_{n=1}^{\infty}\theta_{n}\|x_{n}-x_{n-1}\|^{2}<\infty$, and $\inf_{n\in\mathbb{N}}\alpha_{n}>0$ and $\sup_{n\in\mathbb{N}}\alpha_{n}<1$.
The second condition that $\sum_{n=1}^{\infty}\theta_{n}\|x_{n}-x_{n-1}\|^{2}<\infty$, taken by Mainge [Mainge2008] is often difficult to verify in practice. Boţ et al. [Bot2015] proposed a modification in algorithm (8) by replacing this condition with an alternative criterion that the sequence $\{\theta_{n}\}$ is nondecreasing in $0,\theta]$ with $\theta_{1}=0$ and $\theta\in[0,1)$ such that
$\delta>\frac{\alpha^{2}(1+\alpha)+\alpha\sigma}{1-\alpha^{2}}\text{ and }0<\lambda\leq\lambda_{n}\leq\frac{\delta-\alpha[\alpha(1+\alpha)+\alpha\delta+\sigma]}{\delta[1+\alpha(1+\alpha)+\alpha\delta+\sigma]}\text{ for all }n\in\mathbb{N},$ (9)
where $\alpha,\delta,\sigma>0$ are some suitable parameters.
Recently, Dixit et al. [Dixit2020] proposed the inertial normal S-iteration method for the solution of problem (4) for a nonexpansive operator, which in practice converges faster than the Mann iteration method, inertial Mann iteration method and normal S-iteration method.
The verification of condition (9), as considered by [Bot2015], is a challenging task. To address this issue, Sahu [Sahusoft2020] proposed a new convergence criterion for the inertial Mann iteration. This criterion assumes that the sequences $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ satisfy the following conditions: there exist constants $a,b>0$ such that $0<a\leq\alpha_{n}\leq b<1$ for all $n\in\mathbb{N}$, the sequence $\{\theta_{n}\}$ is nondecreasing in $[0,1)$, and there exists $\theta\in(0,1)$ such that $0\leq\theta_{n}\leq\theta<\frac{\epsilon}{1+\epsilon+\max\{1,\epsilon\}}$ for all $n\in\mathbb{N}$, where $\epsilon=\frac{1-b}{b}$. These conditions are easier to verify in practice.
Motivated by the work of Sahu [Sahusoft2020] and Bergmann et al. [Bergmann2016], the main purpose of this paper is to introduce an inertial Douglas-Rachford method and $p$-accelerated normal S-Douglas-Rachford method for solving problem (2) and study its convergence analysis. Our proposed methods are novel and of interest in the following sense:
  • (i)
    Compared with the classical Douglas–Rachford method on Hadamard manifolds Bergmann et al. [Bergmann2016], we propose two accelerated variants on Hadamard manifolds, namely an inertial Douglas-Rachford method and $p$-accelerated normal S-Douglas-Rachford method. To the best of our knowledge, such inertial and $p$-accelerated Douglas–Rachford schemes have not been previously analyzed in the Riemannian setting.
  • (ii)
    Unlike Sahu [Sahusoft2020] and other works on $p$-accelerated normal S-iteration in Hilbert spaces, we extend the $p$-acceleration idea to Douglas–Rachford splitting methods on Hadamard manifolds. Although some of the conditions are similar to the Hilbert space case, this extension is not straightforward because Hadamard manifolds do not have a linear structure. It requires careful use of geodesic convexity, exponential maps, and comparison geometry.
  • (iii)
    Compared with the inertial Douglas–Rachford and inertial Mann schemes studied in Hilbert spaces by Boţ et al. [Bot2015] and Mainge [Mainge2008], our analysis is carried out entirely in the nonlinear setting of Hadamard manifolds. Moreover, we use different parameter conditions, which are easier to verify in practice than those in Boţ et al. [Bot2015] and Mainge [Mainge2008].
  • (iv)
    We analyze the convergence rates of our proposed algorithms. Specifically, the $p$-accelerated normal S-iteration method converges at a rate of order $o(1/\sqrt{n})$, while the inertial Mann iteration method converges in the minimal residual sense, also at a rate of order $o(1/\sqrt{n})$.
We give an application of these algorithms to minimizing functionals containing multiple summands by introducing a parallel inertial Douglas-Rachford method and parallel $p$-accelerated Douglas-Rachford method. We also apply our parallel Douglas-Rachford methods to solve the generalized Heron problems in Hadamard manifolds and provide numerical experiments.
2 Preliminaries
Let $C$ be a nonempty subset of a metric space $(X,d)$ and let $T\colon C\to C$ be an operator. Then $T$ is said to be nonexpansive if $d(Tx,Ty)\leq d(x,y)\text{ for all }x,y\in C.$ We denote the set of fixed points of an operator $T\colon C\to C$ by $\operatorname*{Fix}(T)$, thus, $\operatorname*{Fix}(T)=\{x\in C\colon Tx=x\}.$
For a sequence $\{x_{n}\}$ in $C$, define $R_{T,\{x_{n}\}}:\mathbb{N}\to[0,\infty)$ by
$R_{T,\{x_{n}\}}(n)=\min_{1\leq i\leq n}d(x_{i},Tx_{i})\text{ for all }n\in\mathbb{N}.$
Observe that $R_{T,\{x_{n}\}}(n)\leq d(x_{i},Tx_{i})$ for all $i=1,2,\ldots,n$ and $\{R_{T,\{x_{n}\}}(n)\}$ is decreasing. We use the notation $o$ to mean that $s_{n}=o(1/t_{n})$ if and only if $\lim_{n\to\infty}s_{n}t_{n}=0.$
Lemma 1 ([Dong2015comments]).
Let $\{a_{n}\}$ and $\{b_{n}\}$ be sequences of positive real numbers such that $\sum_{n=1}^{\infty}a_{n}b_{n}<\infty.$ Suppose that $\sum_{n=1}^{\infty}a_{n}=\infty$ and that the sequence $\{b_{n}\}$ is decreasing. Then $b_{n}=o\!\left(\frac{1}{\sum_{i=1}^{n}a_{i}}\right).$
Definition 2 ([SahuNFAO]).
Let $C$ be a nonempty subset of a metric space $(X,d)$ and let $T\colon C\to C$ be an operator with $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be a sequence in $C$. We say that $\{x_{n}\}$ has
  1. (a)
    the limit existence property (for short, the LE property) for $T$ if $\lim_{n\to\infty}d(x_{n},v)$ exists for all $v\in\operatorname*{Fix}(T)$;
  2. (b)
    the approximate fixed point property (for short, the AF point property) for $T$ if
    $\lim_{n\to\infty}d(x_{n},Tx_{n})=0;$
  3. (c)
    the LEAF point property if $\{x_{n}\}$ satisfies both LE property and AF property.
Lemma 3 ([Sahusoft2020]).
Let $X$ be a Hilbert space and $\alpha\geq 0$. Then,
$\|x-\alpha y\|^{2}\geq(1-\alpha)\|x\|^{2}-\alpha(1-\alpha)\|y\|^{2}\text{ for all }x,y\in X.$
Lemma 4 ([Alvarez2004]).
Let $\{a_{n}\},\{b_{n}\}$ and $\{\theta_{n}\}$ be sequences in $0,\infty)$ such that
$a_{n+1}\leq a_{n}+\theta_{n}(a_{n}-a_{n-1})+b_{n}\text{ for all }n\in\mathbb{N},$
$\sum_{n=1}^{\infty}b_{n}<\infty$ and there exists a real number $\theta$ with $0\leq\theta_{n}\leq\theta<1$ for all $n\in\mathbb{N}$. Then, the following hold:
  1. (i)
    $\sum_{n=1}^{\infty}[a_{n}-a_{n-1}]_{+}<\infty$, where $t]_{+}=\max\{t,0\}$.
  2. (ii)
    There exists $a^{*}\in[0,\infty)$ such that $a_{n}\to a^{*}$.
For the basic definitions and results regarding the geometry of Riemannian manifolds, we refer to [Docarmo, Sakai, Udriste].
Let $\mathcal{M}$ be an $m$-dimensional differentiable manifold and let $x\in\mathcal{M}$. The set of all tangent vectors at the point $x$, denoted by $\mathcal{T}_{x}\mathcal{M}$, is called the tangent space of $\mathcal{M}$ at $x$. The tangent bundle of $\mathcal{M}$ is defined as the disjoint union $\mathcal{T}\mathcal{M}=\bigcup_{x\in\mathcal{M}}\mathcal{T}_{x}\mathcal{M}$. A differentiable manifold $\mathcal{M}$ with a Riemannian metric $\langle\cdot,\cdot\rangle$ is said to be a Riemannian manifold. The induced norm on $\mathcal{T}_{x}\mathcal{M}$ is denoted by $\|\cdot\|_{x}$. The Riemannian distance between $x$ and $y$ in $\mathcal{M}$ is designated by $d(x,y)$. A geodesic joining $x$ to $y$ in the Riemannian manifold $\mathcal{M}$ is said to be a minimal geodesic if its length is equal to $d(x,y)$. We denote geodesic joining $x$ and $y$ by $\gamma(x,y;\cdot)$, i.e., $\gamma(x,y;\cdot)\colon[a,b]\to\mathcal{M}$ is such that, for $a,b\in\mathbb{R}$, we have $\gamma(x,y;a)=x$ and $\gamma(x,y;b)=y$.
Let $\nabla$ be the Riemannian connection associated with the Riemannian manifold $\mathcal{M}$. The parallel transport on the tangent bundle $\mathcal{T}\mathcal{M}$ along the geodesic $\gamma(x,y;\cdot)$ with respect to Riemannian connection $\nabla$ is defined as $P_{\gamma(x,y;b),\gamma(x,y;a)}\colon\mathcal{T}_{\gamma(x,y;a)}\mathcal{M}\to\mathcal{T}_{\gamma(x,y;b)}\mathcal{M}$ such that $P_{\gamma(x,y;b),\gamma(x,y;a)}(v)=V(\gamma(x,y;b))\text{ for all }a,b\in\mathbb{R},v\in\mathcal{T}_{x}\mathcal{M},$ where $V$ is the unique vector field satisfying $\nabla_{\gamma^{\prime}(x,y;t)}V=0$ for all $t\in[a,b]$ and $V(\gamma(x,y;a))=v$. A Riemannian manifold $\mathcal{M}$ is said to be complete if for any $x\in\mathcal{M}$ all geodesic emanating from $x$ are defined for all $t\in\mathbb{R}$. By Hopf-Rinow Theorem [Sakai], if $\mathcal{M}$ is complete, then any pair of points in $\mathcal{M}$ can be joined by a minimal geodesic. Moreover, $(\mathcal{M},d)$ is a complete metric space. A complete simply connected Riemannian manifold of nonpositive sectional curvature is called a Hadamard manifold. Since in a Hadamard manifold $\mathcal{M}$, any two points can be joined by a unique geodesic, we denote the parallel transport by $P_{y,x}$ instead of $P_{\gamma(x,y;b),\gamma(x,y;a)}$. If $\mathcal{M}$ is a Hadamard manifold, then the exponential map $\exp_{x}\colon\mathcal{T}_{x}\mathcal{M}\to\mathcal{M}$ at $x\in\mathcal{M}$ is defined by $\exp_{x}v=\gamma_{x,v}(1)$ for all $v\in\mathcal{T}_{x}\mathcal{M}$, where $\gamma_{x,v}\colon\mathbb{R}\to\mathcal{M}$ is a unique geodesic starting from $x$ with velocity $v$, i.e., $\gamma_{x,v}(0)=x$ and $\gamma^{\prime}_{x,v}(0)=v$. It is known that $\exp_{x}(tv)=\gamma_{x,v}(t)$ for each real number $t$. For the exponential map, the following properties hold:
  1. (i)
    $\exp_{x}\mathbf{0}=\gamma_{x,\mathbf{0}}(1)=\gamma_{x,v}(0)=x$, where $\mathbf{0}$ is the zero tangent vector
  2. (ii)
    Inverse of $\exp_{x}$, $\exp_{x}^{-1}\colon\mathcal{M}\to\mathcal{T}_{x}\mathcal{M}$ is given by $\exp_{x}^{-1}y=v,$ where $\gamma_{x,v}(1)=\exp_{x}v=y$.
  3. (iii)
    $\exp_{x}$ is a diffeomorphism on $\mathcal{T}_{x}\mathcal{M}$ and for any $x,y\in\mathcal{M}$, we have $d(x,y)=\|\exp_{x}^{-1}y\|$.
The following proposition establishes the relationship between geodesics and the exponential map.
Proposition 5 ([Sakai]).
Let $\mathcal{M}$ be a Hadamard manifold and $x\in\mathcal{M}$. Then $\exp_{x}\colon\mathcal{T}_{x}\mathcal{M}\to\mathcal{M}$ is a diffeomorphism and for any two points $x,y\in\mathcal{M}$, there exists a unique unit-speed geodesic joining $x$ to $y$, which is in fact a minimal geodesic. Let $\gamma(x,y;\cdot)\colon[0,1]\to\mathcal{M}$ be the desired geodesic joining $x$ to $y$, i.e., $x=\gamma(x,y;0)$ and $y=\gamma(x,y;1)$. Then
$\gamma(x,y;t)=\exp_{x}(t\exp_{x}^{-1}y)\text{ for all }t\in[0,1].$
Remark 6 ([Martin2]).
Let $\mathcal{M}$ be a Hadamard manifold. If $x,y\in\mathcal{M}$ and $v\in\mathcal{T}_{y}\mathcal{M}$, then
$\langle v,-\exp_{y}^{-1}x\rangle=\langle v,P_{x,y}\exp_{x}^{-1}y\rangle=\langle P_{x,y}v,\exp_{x}^{-1}y\rangle.$
Hadamard manifolds have the nice feature that they resemble convexity properties from Euclidean spaces. A subset $C$ of a Hadamard manifold $\mathcal{M}$ is said to be geodesically convex if, for any two points $x$ and $y$ in $C$, the geodesic joining $x$ to $y$ is contained in $C$. A function $f\colon\mathcal{M}\to\mathbb{R}$ is said to be geodesically convex if, for any geodesic $\gamma(x,y;\cdot)\colon[a,b]\to\mathcal{M}$, the composition function $f\circ\gamma(x,y;\cdot)\colon[a,b]\to\mathbb{R}$ is convex, i.e.,
$f\circ\gamma(x,y;(1-t)a+tb)\leq(1-t)(f\circ\gamma(x,y;a))+t(f\circ\gamma(x,y;b)),$
for any $a,b\in\mathbb{R}$, $x,y\in\mathcal{M}$ and $t\in[0,1]$.
The following proposition highlights the convexity of the distance function.
Proposition 7 ([Sakai]).
Let $\mathcal{M}$ be a Hadamard manifold and $d\colon\mathcal{M}\times\mathcal{M}\to\mathbb{R}$ be the metric function. Then $d$ is a geodesically convex function with respect to product Riemannian metric, i.e., given any pair of geodesics $\gamma(x_{1},y_{1};\cdot)\colon[0,1]\to\mathcal{M}$ and $\gamma(x_{2},y_{2};\cdot)\colon[0,1]\to\mathcal{M}$, for some $x_{1},y_{1},x_{2},y_{2}\in\mathcal{M}$, the following hold:
  1. (i)
    $d(\gamma(x_{1},y_{1};t),\gamma(x_{2},y_{2};t))\leq(1-t)d(x_{1},x_{2})+td(y_{1},y_{2})\text{ for all }t\in[0,\,1].$
  2. (ii)
    $d(\gamma(x_{1},y_{1};t),z)\leq(1-t)d(x_{1},z)+td(y_{1},z)\text{ for all }t\in[0,1]\text{ and }z\in\mathcal{M}.$
Proposition 8 ([ShikherCNSNS2024]).
Let $\mathcal{M}$ be a Hadamard manifold and $x,y,z\in\mathcal{M}$. Let $u\in\mathcal{M}$ be a point on the geodesic joining $x,y$, i.e., $u=\gamma(x,y;t)$, for some $t\in[0,1]$, then
$d^{2}(u,z)\leq(1-t)d^{2}(x,z)+td^{2}(y,z)-t(1-t)d^{2}(x,y).$
Proposition 9.
Let $\mathcal{M}$ be a Hadamard manifold and $x,y\in\mathcal{M}$. Then, for $z\in\mathcal{M}$ and $\theta\in[0,1]$, the following inequality holds:
$d^{2}(\exp_{x}(-\theta\exp_{x}^{-1}y),z)\leq(1+\theta)d^{2}(x,z)-\theta d^{2}(y,z)+\theta(1+\theta)d^{2}(x,y).$
(The proof is given in Appendix A.)
Let $\mathcal{M}$ be a Hadamard manifold and let $N\in\mathbb{N}$. Then, the product space of manifold $\mathcal{M}$ is defined by $\mathcal{M}^{N}=\{x=(x_{1},\ldots,x_{N})\colon x_{i}\in\mathcal{M}\text{ for }i=1,2,\ldots,N\}.$ Note that product space $\mathcal{M}^{N}$ is again a Hadamard manifold. The tangent space of $\mathcal{M}^{N}$ at $x=(x_{1},\ldots,x_{N})$ is $\mathcal{T}_{x}\mathcal{M}^{N}=\mathcal{T}_{x_{1}}\mathcal{M}\times\cdots\times\mathcal{T}_{x_{N}}\mathcal{M}$. The distance between two points $x=(x_{1},\ldots,x_{N})$ and $y=(y_{1},\ldots,y_{N})$ in $\mathcal{M}^{N}$ is defined by
$\rho(x,y)=\left(\sum_{i=1}^{N}d^{2}(x_{i},y_{i})\right)^{1/2}.$
Moreover, the geodesic in $\mathcal{M}^{N}$ joining $x=(x_{1},\ldots,x_{N})$ to $y=(y_{1},\ldots,y_{N})$ is given componentwise by $\gamma(x,y;t)=\bigl(\gamma(x_{i},y_{i};t),\ldots,\gamma(x_{N},y_{N};t)\bigr)\text{ for all }t\in[0,1]$.
We now recall an important property related to the convergence of sequences satisfying the LEAF point (limit existence and approximate fixed point) property.
Proposition 10 ([ShikherCNSNS2024]).
Let $C$ be a nonempty and closed subset of a Hadamard manifold $\mathcal{M}$ and let $T\colon C\to C$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be a sequence in $C$ satisfying the LEAF point property for $T$. Then, $\{x_{n}\}$ converges to a fixed point of $T$.
[…]
$\operatorname{Prox}_{\!\lambda f}(x)=\operatorname*{argmin}_{y\in\mathcal{M}}\left(f(y)+\frac{1}{2\lambda}d^{2}(x,y)\right)\text{ for all }x\in\mathcal{M}.$
The proximal operator of a proper, geodesically convex and lower semicontinuous function is firmly nonexpansive operator (see [Jost1995, Martin3]), i.e., for any $x,y\in\mathcal{M}$ and $t\in[0,1]$
$d(\operatorname{Prox}_{\!\lambda f}(x),\operatorname{Prox}_{\!\lambda f}(y))\leq d(\gamma(x,\operatorname{Prox}_{\!\lambda f}(x);t),\gamma(y,\operatorname{Prox}_{\!\lambda f}(y);t)).$ (10)
[…]
$R_{\lambda f}(x)=\exp_{\operatorname*{Prox}_{\lambda f}(x)}(-\exp_{\operatorname*{Prox}_{\lambda f}(x)}^{-1}(x))\text{ for all }x\in\mathcal{M}.$
That is, $R_{\lambda f}(x)$ is the reflection of $x$ with respect to $\operatorname*{Prox}_{\lambda f}(x)$ along the unique geodesic joining $x$ to $\operatorname*{Prox}_{\lambda f}(x)$.
Lemma 11 ([Bergmann2016]).
Let $\mathcal{M}$ be an Hadamard manifold, $\lambda>0$ and $c\in\mathcal{M}$. Define $f\colon\mathcal{M}\to\mathbb{R}$ by $f(x)=d(c,x)\text{ for all }x\in\mathcal{M}$. Then, $\operatorname*{Prox}_{\lambda f}(x)=\gamma(x,c,t)$, where $t=\min\big\{\tfrac{\lambda}{d(x,c)},1\big\}$.
[…]
Theorem 12 ([Bergmann2016]).
[…]
$\iota_{C}=\begin{cases}0,&\text{ if }x\in C,\\ \infty,&\text{otherwise}.\end{cases}$
Next, we present some results for functions with nonexpansive reflections.
Proposition 13 ([Leon2013, Bergmann2016]).
Let $C$ be a nonempty, closed and geodesically convex subset of Hadamard manifold $\mathcal{M}$ with constant sectional curvature. Then
$d(R_{\iota_{C}}(x),R_{\iota_{C}}(y))\leq d(x,y)\text{ for all }x,y\in C.$
Theorem 14 ([Bergmann2016]).
Let $\mathcal{M}$ be a Hadamard Manifold, $\lambda>0$ and $c\in\mathcal{M}$. Let $g\colon\mathcal{M}\to\mathbb{R}$ be defined by $g(x)=d(c,x)\text{ for all }x\in\mathcal{M}$. Then, the reflection $R_{\lambda g}$ is nonexpansive.
Proposition 15.
[…]
(The proof is given in Appendix B.)
3 Efficient Douglas-Rachford Algorithms on Hadamard Manifolds
In this section, we introduce two algorithms for finding fixed points of nonexpansive operators on Hadamard manifolds. We also discuss the consequences of these algorithms for solving the minimization problem (2).
First, we introduce a method that incorporates an inertial extrapolation acceleration term, given as follows:
Algorithm 16.
Let $\mathcal{M}$ be a Hadamard manifold and $T\colon\mathcal{M}\to\mathcal{M}$ be an operator.
Initialization: Choose $x_{0},x_{1}\in\mathcal{M}$ arbitrary.
Iterative step: For the current $n^{th}$ iterate, compute the $(n+1)^{th}$ iteration as follows:
$\begin{cases}y_{n}=\exp_{x_{n}}(-\theta_{n}\exp_{x_{n}}^{-1}x_{n-1}),\\ x_{n+1}=\gamma(y_{n},Ty_{n};\alpha_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (11)
where $\{\alpha_{n}\}$ is a sequence in (0,1)(0,1)and $\{\theta_{n}\}$ is a sequence in $[0,1)$.
Note that method (11) is an extension of the inertial Mann iteration method (8) from Hilbert spaces to the setting of Hadamard manifolds. Therefore, we also refer to it here as the inertial Mann iteration method.
For the convergence analysis of inertial Mann method (11), we assume that the sequences $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ satisfy the following conditions:
  1. (C1)
    $0<a\leq\alpha_{n}\leq b<1$ for all $n\in\mathbb{N}$ for some $a,b>0$,
  2. (C2)
    $\{\theta_{n}\}$ is a nondecreasing sequence in $[0,1)$,
  3. (C3)
    There exists $\theta\in(0,1)$ such that $0\leq\theta_{n}\leq\theta<\frac{\epsilon}{1+\epsilon+\max\{1,\epsilon\}}$ for all $n\in\mathbb{N}$, where $\epsilon=\frac{1-b}{b}$.
Remark 17.
The upper bound for the inertial parameter $\{\theta_{n}\}$ in the assumption of Boţ et al. [Bot2015] (see (9)) depends on several auxiliary parameters $\alpha$, $\delta$, and $\sigma$. Consequently, determining an admissible range for $\{\theta_{n}\}$ requires careful tuning of these coupled parameters and is not straightforward. In contrast, under our condition (C3), an explicit upper bound for $\{\theta_{n}\}$ can be directly obtained once an upper bound for the sequence $\{\alpha_{n}\}$ is fixed. Intuitively, the parameter $\theta_{n}$ controls the amount of inertia introduced at each iteration, while the bound involving $\epsilon=(1-b)/b$ ensures that this inertial effect remains sufficiently small relative to the relaxation parameter $\alpha_{n}$. This allows the parameters to be chosen independently, making the conditions easier to verify and more convenient to implement in practice than those in [Bot2015].
Theorem 18.
Let $\mathcal{M}$ be a Hadamard manifold and let $T\colon\mathcal{M}\to\mathcal{M}$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be a sequence generated by inertial Mann method (11), where $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ are sequences satisfying the conditions (C1), (C2) and (C3). For $v\in\operatorname*{Fix}(T)$, define $\phi_{n}=d^{2}(x_{n},v),\psi_{n}=\phi_{n}-\theta_{n}\phi_{n-1}+K_{n}d^{2}(x_{n},x_{n-1})$, where $K_{n}=\theta_{n}(1+\theta_{n}+\epsilon(1-\theta_{n}))$. Assume that $\psi_{1}>0$. Then, we have the following:
  • (i)
    $\{\psi_{n}\}$ is nonincreasing.
  • (ii)
    $\sum_{n=1}^{\infty}d^{2}(x_{n+1},x_{n})\leq\frac{\psi_{1}}{K(1-\theta)},$ where $K=\epsilon-\theta(1+\epsilon+\max\{1,\epsilon\})>0$.
  • (iii)
    $\lim_{n\to\infty}d(x_{n},v)$ and $\lim_{n\to\infty}d(x_{n},v)$ exists and $\lim_{n\to\infty}d(y_{n},Ty_{n})=0$.
  • (iv)
    $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$.
Proof.
(i) Let $v\in\operatorname*{Fix}(T)$. From (11) and Proposition 9, we have
$d^{2}(y_{n},v)\leq(1+\theta_{n})d^{2}(x_{n},v)-\theta_{n}d^{2}(x_{n-1},v)+\theta_{n}(1+\theta_{n})d^{2}(x_{n},x_{n-1}).$ (12)
From (11) and Proposition 8, we have
$d^{2}(x_{n+1},v)$ $=(1-\alpha_{n})d^{2}(y_{n},v)+\alpha_{n}d^{2}(Ty_{n},v)-\alpha_{n}(1-\alpha_{n})d^{2}(y_{n},Ty_{n})$
$\leq(1-\alpha_{n})d^{2}(y_{n},v)+\alpha_{n}d^{2}(y_{n},v)-\alpha_{n}(1-\alpha_{n})d^{2}(y_{n},Ty_{n})$
$=d^{2}(y_{n},v)-\alpha_{n}(1-\alpha_{n})d^{2}(y_{n},Ty_{n}).$ (13)
Consider the geodesic triangle $\triangle(x_{n},x_{n-1},v)$ and its comparison triangle $\triangle(x_{n}^{\prime},x_{n-1}^{\prime},v^{\prime})$. Then
$d(x_{n},x_{n-1})=\|x_{n}^{\prime}-x_{n-1}^{\prime}\|,d(x_{n-1},v)=\|x_{n-1}^{\prime}-v^{\prime}\|\text{ and }d(x_{n},v)=\|x_{n}^{\prime}-v^{\prime}\|.$ (14)
Since $y_{n}=\exp_{x_{n}}(-\theta_{n}\exp_{x_{n}}^{-1}x_{n-1})$, its comparison point will be
$y_{n}^{\prime}=(1+\theta_{n})x_{n}^{\prime}+(-\theta_{n})x_{n-1}^{\prime}.$ (15)
Now consider the geodesic triangle $\triangle(y_{n},{Ty_{n}},v)$. Without loss of generality, we can choose its comparison triangle to be $\triangle(y_{n}^{\prime},Ty_{n}^{\prime},\overline{v})$, where $y_{n}^{\prime}$ is the same as defined in (15). Since $x_{n+1}=\exp_{y_{n}}(\alpha_{n}\exp_{y_{n}}^{-1}Ty_{n})$, its comparison point will be
$x_{n+1}^{\prime}=(1-\alpha_{n})y_{n}^{\prime}+\alpha_{n}Ty_{n}^{\prime}.$ (16)
Without loss of generality, we can also choose $Ty_{n}^{\prime}$ such that
$d(x_{n},x_{n+1})=\|x_{n}^{\prime}-x_{n+1}^{\prime}\|$ (17)
as shown in Figure 1. Then
$d(y_{n},Ty_{n})=\|y_{n}^{\prime}-Ty_{n}^{\prime}\|,d(Ty_{n},v)=\|Ty_{n}^{\prime}-\overline{v}\|\text{ and }d(y_{n},v)=\|y_{n}^{\prime}-\overline{v}\|.$ (18)
From (16), we have $x_{n+1}^{\prime}=\alpha_{n}(Ty_{n}^{\prime}-y_{n}^{\prime})+y_{n}^{\prime}$, which implies that
$Ty_{n}^{\prime}-y_{n}^{\prime}=\frac{1}{\alpha_{n}}(x_{n+1}^{\prime}-y_{n}^{\prime}).$ (19)
Since $x_{n+1}^{\prime}$ is comparison point of $x_{n+1}$, we have $d(x_{n+1},y_{n})=\|x_{n+1}^{\prime}-y_{n}^{\prime}\|$. It follows from (18) and (19) that
$d^{2}(Ty_{n},y_{n})=\|Ty_{n}^{\prime}-y_{n}^{\prime}\|^{2}=\frac{1}{\alpha_{n}^{2}}\|x_{n+1}^{\prime}-y_{n}^{\prime}\|^{2}=\frac{1}{\alpha_{n}^{2}}d^{2}(x_{n+1},y_{n}).$
Thus, from (13), we have
$d^{2}(x_{n+1},v)$ $\leq d^{2}(y_{n},v)-\frac{1-\alpha_{n}}{\alpha_{n}}d^{2}(x_{n+1},y_{n})\leq d^{2}(y_{n},v)-\frac{1-b}{b}d^{2}(x_{n+1},y_{n}).$
It follows from (12) that
$d^{2}(x_{n+1},v)\leq(1+\theta_{n})d^{2}(x_{n},v)-\theta_{n}d^{2}(x_{n-1},v)+\theta_{n}(1+\theta_{n})d^{2}(x_{n},x_{n-1})-\epsilon d^{2}(x_{n+1},y_{n}).$ (20)
By (15) and Lemma 3, we have
$d^{2}(x_{n+1},y_{n})$ $=\|x_{n+1}^{\prime}-y_{n}^{\prime}\|^{2}=\|x_{n+1}^{\prime}-((1+\theta_{n})x_{n}^{\prime}+(-\theta_{n})x_{n-1}^{\prime})\|^{2}$
$=\|(x_{n+1}^{\prime}-x_{n}^{\prime})-\theta_{n}(x_{n}^{\prime}-x_{n-1}^{\prime})\|^{2}$
$\geq(1-\theta_{n})\|x_{n+1}^{\prime}-x_{n}^{\prime}\|^{2}-\theta_{n}(1-\theta_{n})\|x_{n}^{\prime}-x_{n-1}^{\prime}\|^{2}.$
using (14) and (17), we have $d(x_{n},x_{n-1})=\|x_{n}^{\prime}-x_{n-1}^{\prime}\|$ and $d(x_{n},x_{n+1})=\|x_{n}^{\prime}-x_{n+1}^{\prime}\|$. Thus
$d^{2}(x_{n+1},y_{n})\geq(1-\theta_{n})d^{2}(x_{n+1},x_{n})-\theta_{n}(1-\theta_{n})d^{2}(x_{n},x_{n-1}).$ (21)
$x_{n-1}^{\prime}$$v^{\prime}$$x_{n}^{\prime}$$y_{n}^{\prime}$$Ty_{n}^{\prime}$$\bar{v}$$x_{n+1}^{\prime}$$d(x_{n},x_{n+1})$
Figure 1: Selection of comparison points
From (20) and (21), we get
$d^{2}(x_{n+1},v)\leq$ $(1+\theta_{n})d^{2}(x_{n},v)-\theta_{n}d^{2}(x_{n-1},v)+\theta_{n}(1+\theta_{n})d^{2}(x_{n},x_{n-1})$
$-\epsilon[(1-\theta_{n})d^{2}(x_{n},x_{n+1})-\theta_{n}(1-\theta_{n})d^{2}(x_{n},x_{n-1})]$
$=$ $(1+\theta_{n})d^{2}(x_{n},v)-\theta_{n}d^{2}(x_{n-1},v)-\epsilon(1-\theta_{n})d^{2}(x_{n+1},x_{n})$
$+\theta_{n}(1+\theta_{n}+\epsilon(1-\theta_{n}))d^{2}(x_{n},x_{n-1}).$ (22)
Take $\phi_{n}=d^{2}(x_{n},v),\psi_{n}=\phi_{n}-\theta_{n}\phi_{n-1}+K_{n}d^{2}(x_{n},x_{n-1})$, where $K_{n}=\theta_{n}(1+\theta_{n}+\epsilon(1-\theta_{n}))$. Then, from (22), we have
$\phi_{n+1}\leq(1+\theta_{n})\phi_{n}-\theta_{n}\phi_{n-1}-\epsilon(1-\theta_{n})d^{2}(x_{n+1},x_{n})+K_{n}d^{2}(x_{n},x_{n-1})\text{ for all }n\in\mathbb{N}.$
By the definition of $\psi_{n}$, we obtain
$\psi_{n+1}-\psi_{n}$
$=\phi_{n+1}-\theta_{n+1}\phi_{n}+K_{n+1}d^{2}(x_{n+1},x_{n})-\phi_{n}+\theta_{n}\phi_{n-1}-K_{n}d^{2}(x_{n},x_{n-1})$
$=\phi_{n+1}-(1+\theta_{n+1})\phi_{n}+\theta_{n}\phi_{n-1}+K_{n+1}d^{2}(x_{n+1},x_{n})-K_{n}d^{2}(x_{n},x_{n-1})$
$\leq\phi_{n+1}-(1+\theta_{n})\phi_{n}+\theta_{n}\phi_{n-1}+K_{n+1}d^{2}(x_{n+1},x_{n})-K_{n}d^{2}(x_{n},x_{n-1})$
$\leq-\epsilon(1-\theta_{n})d^{2}(x_{n+1},x_{n})+K_{n}d^{2}(x_{n},x_{n-1})+K_{n+1}d^{2}(x_{n+1},x_{n})-K_{n}d^{2}(x_{n},x_{n-1})$
$=-(\epsilon(1-\theta_{n})-K_{n+1})d^{2}(x_{n+1},x_{n})\text{ for all }n\in\mathbb{N}.$ (23)
Note that
$K_{n}=\theta_{n}(1+\theta_{n}+\epsilon(1-\theta_{n}))\leq\theta_{n}(1+\max\{1,\epsilon\})\text{ for all }n\in\mathbb{N}$
and hence
$-[\epsilon(1-\theta_{n})-K_{n+1}]$ $\leq-\epsilon+\epsilon\theta_{n}+\theta_{n+1}(1+\max\{1,\epsilon\})\leq-\epsilon+\theta(1+\epsilon+\max\{1,\epsilon\}).$
From condition (C3), we see that $K=\epsilon-\theta(1+\epsilon+\max\{1,\epsilon\})>0$. Using (23), we obtain
$\psi_{n+1}-\psi_{n}\leq-Kd^{2}(x_{n+1},x_{n})\text{ for all }n\in\mathbb{N}.$ (24)
Therefore $\{\psi_{n}\}$ is nonincreasing.
(ii) Note $\psi_{1}=\phi_{1}-\theta_{1}\phi_{0}+K_{1}d^{2}(x_{1},x_{0})>0$. Since $\{\theta_{n}\}$ is bounded above by $\theta$, we get
$\psi_{n}=\phi_{n}-\theta_{n}\phi_{n-1}+K_{n}d^{2}(x_{n},x_{n-1})\geq\phi_{n}-\theta_{n}\phi_{n-1}\geq\phi_{n}-\theta\phi_{n-1}.$
Since $\{\psi_{n}\}$ is nonincreasing, we get
$\phi_{n}-\theta\phi_{n-1}\leq\psi_{n}\leq\psi_{1}\text{ for all }n\in\mathbb{N}.$ (25)
Thus,
$\phi_{n}\leq\theta\phi_{n-1}+\psi_{1}\leq\theta(\theta\phi_{n-2}+\psi_{1})+\psi_{1}\leq\cdots$ $\leq\theta^{n}\phi_{0}+\psi_{1}\sum_{k=0}^{n-1}\theta^{k}$
$\leq\theta^{n}\phi_{0}+\frac{\psi_{1}}{1-\theta}\text{ for all }n\in\mathbb{N}.$
It follows from (24) and (25) that
$K\sum_{k=1}^{n}d^{2}(x_{k+1},x_{k})\leq\psi_{1}-\psi_{n+1}\leq\psi_{1}+\theta\phi_{n}\leq\psi_{1}+\theta\left(\phi_{0}\theta^{n}+\frac{\psi_{1}}{1-\theta}\right)=\theta^{n+1}\phi_{0}+\frac{\psi_{1}}{1-\theta}.$
Since $\theta^{n}\to 0$ as $n\to\infty$, we have
$\sum_{n=1}^{\infty}d^{2}(x_{n+1},x_{n})\leq\frac{\psi_{1}}{K(1-\theta)}.$ (26)
(iii) It follows from (26) that
$\lim_{n\to\infty}d(x_{n},x_{n+1})=0.$ (27)
Using (14), (17) and (27), we deduce that
$d(x_{n+1},y_{n})$ $=\|x_{n+1}^{\prime}-y_{n}^{\prime}\|=\|x_{n+1}^{\prime}-[(1+\theta_{n})x_{n}^{\prime}+(-\theta_{n}){x_{n-1}^{\prime}}]\|$
$=\|x_{n+1}^{\prime}-x_{n}^{\prime}+\theta_{n}(x_{n-1}^{\prime}-x_{n}^{\prime})\|\leq\|x_{n+1}^{\prime}-x_{n}^{\prime}\|+\theta_{n}\|{x_{n-1}^{\prime}}-x_{n}^{\prime}\|$
$=d(x_{n+1},x_{n})+\theta_{n}d(x_{n-1},x_{n})\to 0\text{ as }n\to\infty.$ (28)
Observe from (27) and (28) that
$d(x_{n},y_{n})\leq d(x_{n},x_{n+1})+d(x_{n+1},y_{n})\to 0\text{ as }n\to\infty.$ (29)
From (22) and (24), we have
$\phi_{n+1}\leq(1+\theta_{n})\phi_{n}-\theta_{n}\phi_{n-1}+\theta(1+\max\{1,\epsilon\})d^{2}(x_{n},x_{n-1})\text{ for all }n\in\mathbb{N}.$
From (26) and Lemma 4, we obtain that $\lim_{n\to\infty}d(x_{n},v)$ exists. We may assume that $\lim_{n\to\infty}d(x_{n},v)=l>0$. It follows from (29) that $\lim_{n\to\infty}d(y_{n},v)=l.$ Hence from (13), we obtain that $\lim_{n\to\infty}d(y_{n},Ty_{n})=0.$
(iv) Note that $\lim_{n\to\infty}d(y_{n},v)=l$ exists and $\lim_{n\to\infty}d(y_{n},Ty_{n})=0$. Hence $\{y_{n}\}$ satisfies the LEAF point property, therefore by Proposition 10, $\{y_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$. Since $\lim_{n\to\infty}d(x_{n},y_{n})=0$, we have $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$. ∎
Remark 19.
Theorem 18 is the extension of Theorem 3.3 of Sahu [Sahusoft2020] from Hilbert spaces to the setting of Hadamard manifolds.
Remark 20.
Note that the sequence $\{d(x_{n},Tx_{n})\}$ may not be decreasing for the inertial Mann iteration method (11), even when $T$ is a nonexpansive operator. In such a situation, the accuracy of the sequence $\{x_{n}\}$ in approximating a fixed point of $T$ cannot be measured by $\{d(x_{n},Tx_{n})\}$. Therefore, we use $R_{T,\{x_{n}\}}(n)$ to evaluate the accuracy of $\{x_{n}\}$ in approximating a fixed point of $T$ when $\{d(x_{n},Tx_{n})\}$ is not essentially decreasing.
In the next result, we study the rate of convergence of the inertial Mann iteration method (11) in terms of $R_{T,\{x_{n}\}}(n)$.
Theorem 21.
Let $\mathcal{M}$ be a Hadamard manifold and let $T\colon\mathcal{M}\to\mathcal{M}$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be a sequence generated by inertial Mann method (11) with $x_{0}=x_{1}\in\mathcal{M}$, where $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ are sequences satisfying the conditions (C1), (C2) and (C3). Then $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$ with the following rate of convergence:
$R_{T,\{x_{n}\}}(n)=o\left(\frac{1}{\sqrt{n}}\right).$
In fact for any $v\in\operatorname*{Fix}(T)$
$(R_{T,\{x_{n}\}}(n))^{2}\leq\frac{5}{a(1-b)}\left(1+\frac{\theta^{2}(1+\theta)}{K(1-\theta)^{2}}+\frac{\theta(1+\theta)}{K(1-\theta)}\right)\frac{d^{2}(x_{0},v)}{n},$
where $K=\epsilon-\theta(1+\epsilon+\max\{1,\epsilon\})$.
Proof.
Let $v\in\operatorname*{Fix}(T)$. From (12) and (13), we have
$d^{2}(x_{n+1},v)$ $\leq(1+\theta_{n})d^{2}(x_{n},v)-\theta_{n}d^{2}(x_{n-1},v)+\theta_{n}(1+\theta_{n})d^{2}(x_{n},x_{n-1})$
$\quad-\alpha_{n}(1-\alpha_{n})d^{2}(y_{n},Ty_{n}).$
It follows from condition (C1) that
$d^{2}(x_{n+1},v)-(1+\theta_{n})d^{2}(x_{n},v)+\theta_{n}d^{2}(x_{n-1},v)$
$\leq-\alpha_{n}(1-\alpha_{n})d^{2}(y_{n},Ty_{n})+\theta_{n}(1+\theta_{n})d^{2}(x_{n},x_{n-1})$
$\leq-a(1-b)d^{2}(y_{n},Ty_{n})+\theta_{n}(1+\theta_{n})d^{2}(x_{n},x_{n-1}),$
which implies that
$(d^{2}(x_{n+1},v)-d^{2}(x_{n},v))-\theta_{n}(d^{2}(x_{n},v)-d^{2}(x_{n-1},v))$
$\leq-a(1-b)d^{2}(y_{n},Ty_{n})+\theta_{n}(1+\theta_{n})d^{2}(x_{n},x_{n-1}).$ (30)
Let $\delta_{n}=\theta_{n}(1+\theta_{n})d^{2}(x_{n},x_{n-1})$, $\phi_{n}=d^{2}(x_{n},v)$, $V_{n}=\phi_{n}-\phi_{n-1}$ and $V_{n}]_{+}=\max\{V_{n},0\}$ for all $n\in\mathbb{N}$. From (30), we have
$a(1-b)d^{2}(y_{n},Ty_{n})$ $\leq\phi_{n}-\phi_{n+1}+\theta_{n}(\phi_{n}-\phi_{n-1})+\delta_{n}$
$=\phi_{n}-\phi_{n+1}+\theta_{n}V_{n}+\delta_{n}.$
$\leq\phi_{n}-\phi_{n+1}+\theta_{n}[V_{n}]_{+}+\delta_{n}.$ (31)
From (26), we have
$\sum_{n=1}^{\infty}d^{2}(x_{n+1},x_{n})\leq\frac{\psi_{1}}{K(1-\theta)}.$ (32)
Note that $x_{0}=x_{1}$. Then $\psi_{1}=d^{2}(x_{1},v)-\theta_{1}d^{2}(x_{0},v)+K_{1}d^{2}(x_{0},x_{1})\leq d^{2}(x_{0},v)$. From (32), we have
$\sum_{n=1}^{\infty}d^{2}(x_{n+1},x_{n})\leq\frac{d^{2}(x_{0},v)}{K(1-\theta)}.$
It follows from condition (C3) that
$\sum_{n=1}^{\infty}\delta_{n}=\sum_{n=1}^{\infty}\theta_{n}(1+\theta_{n})d^{2}(x_{n},x_{n-1})\leq\sum_{n=1}^{\infty}\theta(1+\theta)d^{2}(x_{n},x_{n-1})\leq\frac{\theta(1+\theta)d^{2}(x_{0},v)}{K(1-\theta)}.$ (33)
From (30), we have
$V_{n+1}\leq\theta_{n}V_{n}+\delta_{n}\leq\theta[V_{n}]_{+}+\delta_{n}.$
Therefore
$[V_{n+1}]_{+}\leq\theta[V_{n}]_{+}+\delta_{n}.$
Applying the recurrence repeatedly, we get
$[V_{n+1}]_{+}\leq\theta^{n}[V_{1}]_{+}+\sum_{j=1}^{n}\theta^{j-1}\delta_{n+1-j}.$ (34)
Since $x_{0}=x_{1}$, we get $V_{1}=[V_{1}]_{+}=0$ and $\delta_{1}=0.$ From (34), we have
$\sum_{n=2}^{\infty}[V_{n}]_{+}\leq\frac{1}{1-\theta}\sum_{n=1}^{\infty}\delta_{n}=\frac{1}{1-\theta}\sum_{n=2}^{\infty}\delta_{n}\leq\frac{\theta(1+\theta)d^{2}(x_{0},v)}{K(1-\theta)^{2}}.$ (35)
Since $\phi_{1}=\phi_{0}=d^{2}(x_{0},v)$. From (31), (33) and (35), we get
$a(1-b)\sum_{i=1}^{n}d^{2}(y_{n},Ty_{n})$ $\leq\phi_{1}-\phi_{n}+\theta\sum_{i=1}^{n}[V]_{+}+\sum_{i=2}^{n}\delta_{i}$
$\leq d^{2}(x_{0},v)+\frac{\theta^{2}(1+\theta)d^{2}(x_{0},v)}{K(1-\theta)^{2}}+\frac{\theta(1+\theta)d^{2}(x_{0},v)}{K(1-\theta)}$
$\leq\left(1+\frac{\theta^{2}(1+\theta)}{K(1-\theta)^{2}}+\frac{\theta(1+\theta)}{K(1-\theta)}\right)d^{2}(x_{0},v).$
This implies that
$\sum_{i=1}^{n}d^{2}(y_{n},Ty_{n})\leq\frac{1}{a(1-b)}\left(1+\frac{\theta^{2}(1+\theta)}{K(1-\theta)^{2}}+\frac{\theta(1+\theta)}{K(1-\theta)}\right)d^{2}(x_{0},v).$ (36)
By using the nonexpansivity of $T$, (11) and Proposition 8, we deduce that
$d^{2}(x_{n+1},Tx_{n+1})$ $=d^{2}(\gamma(y_{n},Ty_{n};\alpha_{n}),Tx_{n+1})$
$\leq(1-\alpha_{n})d^{2}(y_{n},Tx_{n+1})+\alpha_{n}d^{2}(Ty_{n},Tx_{n+1})-\alpha_{n}(1-\alpha_{n})d^{2}(y_{n},Ty_{n})$
$\leq(d(y_{n},Ty_{n})+d(Ty_{n},Tx_{n+1}))^{2}+d^{2}(Ty_{n},Tx_{n+1})$
$\leq 2(d^{2}(y_{n},Ty_{n})+d^{2}(Ty_{n},Tx_{n+1}))+d^{2}(Ty_{n},Tx_{n+1})$
$\leq 2d^{2}(y_{n},Ty_{n})+3d^{2}(y_{n},x_{n+1})$
$\leq 2d^{2}(y_{n},Ty_{n})+3d^{2}(y_{n},\gamma(y_{n},Ty_{n};\alpha_{n}))$
$\leq 2d^{2}(y_{n},Ty_{n})+3\alpha_{n}^{2}d^{2}(y_{n},Ty_{n})$
$\leq 5d^{2}(y_{n},Ty_{n}).$
From (36), we have
$\sum_{i=1}^{n}d^{2}(x_{i+1},Tx_{i+1})\leq 5\sum_{i=1}^{n}d^{2}(y_{i},Ty_{i})\leq\frac{5}{a(1-b)}\left(1+\frac{\theta^{2}(1+\theta)}{K(1-\theta)^{2}}+\frac{\theta(1+\theta)}{K(1-\theta)}\right)d^{2}(x_{0},v),$ (37)
which implies that $\sum_{i=1}^{n}(R_{T,\{x_{n}\}}(n))^{2}<\infty$. It follows from Lemma 1 that
$R_{T,\{x_{n}\}}(n)=o\left(\frac{1}{\sqrt{n}}\right).$
Moreover, by (37), we obtain
$(R_{T,\{x_{n}\}}(n))^{2}=\min_{1\leq i\leq n}d^{2}(x_{n+1},Tx_{n+1})\leq\frac{5}{a(1-b)}\left(1+\frac{\theta^{2}(1+\theta)}{K(1-\theta)^{2}}+\frac{\theta(1+\theta)}{K(1-\theta)}\right)\frac{d^{2}(x_{0},v)}{n}.$
Next, we introduce another accelerated method that does not involve an inertial term. Unlike the inertial Mann iteration method, which requires working in the entire space, the following method can be applied on a convex domain.
Algorithm 22.
Let $C$ be a nonempty, closed and geodesically convex subset of a Hadamard manifold $\mathcal{M}$, $p\in\mathbb{N}$ and let $T\colon C\to C$ be an operator.
Initialization: Choose $x_{1}\in C$ arbitrary.
Iterative step: For the current $n^{th}$ iterate, compute the $(n+1)^{th}$ iteration as follows:
$\begin{cases}y_{n}=\gamma(x_{n},Tx_{n};\alpha_{n}),\\ x_{n+1}=T^{p}(y_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (38)
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$.
Note that this method is inspired from the $p$-accelerated normal S-iteration method (7); accordingly, we refer to it by the same name. We now study the convergence of the $p$-accelerated normal S-iteration method (38).
Theorem 23.
Let $C$ be a nonempty, closed and geodesically convex subset of a Hadamard manifold $\mathcal{M}$ and let $T\colon C\to C$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be the sequence generated by $p$-accelerated normal S-iteration method (38), where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ satisfying the condition $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$. Then, we have the following:
  1. (i)
    $\lim_{n\to\infty}d(x_{n},v)$ exists for any $v\in\operatorname*{Fix}(T)$.
  2. (ii)
    $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})<\infty$.
  3. (iii)
    $\{d(x_{n},Tx_{n})\}$ is decreasing.
  4. (iv)
    $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$.
Proof.
(i) Let $v\in\operatorname*{Fix}(T)$. From (38) and Proposition 7, we have
$d(y_{n},v)$ $=d(\gamma(x_{n},Tx_{n};\alpha_{n}),v)$
$\leq(1-\alpha_{n})d(x_{n},v)+\alpha_{n}d(Tx_{n},v)$
$\leq(1-\alpha_{n})d(x_{n},v)+\alpha_{n}d(x_{n},v)=d(x_{n},v).$ (39)
From (38), (39) and nonexpansivity of $T$, we have
$d(x_{n+1},v)=d(T^{p}(y_{n}),v)\leq d(T^{p-1}y_{n},v)\leq\cdots\leq d(y_{n},v)\leq d(x_{n},v).$
(ii) From Proposition 8, we have
$d^{2}(x_{n+1},v)$ $=d^{2}(T^{p}(y_{n}),v)$
$\leq d^{2}(y_{n},v)$
$=d^{2}(\gamma(x_{n},Tx_{n};\alpha_{n}),v)$
$\leq(1-\alpha_{n})d^{2}(x_{n},v)+\alpha_{n}d^{2}(Tx_{n},v)-\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})$
$=d^{2}(x_{n},v)-\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n}).$
It follows that
$\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})\leq d^{2}(x_{n},v)-d^{2}(x_{n+1},v).$
Summing up the above inequality from $n=1$ to $k$, we have
$\sum_{n=1}^{k}\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})\leq\sum_{n=1}^{k}(d^{2}(x_{n},v)-d^{2}(x_{n+1},v))=d^{2}(x_{0},v),$ (40)
which implies that
$\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})<\infty.$ (41)
(iii) Using the nonexpansivity of $T$ and Proposition 7, we have
$d(x_{n+1},Tx_{n+1})$ $=d(T^{p}(y_{n}),T^{p+1}(y_{n}))$
$\leq d(y_{n},Ty_{n})$
$\leq d(y_{n},Tx_{n})+d(Tx_{n},Ty_{n})$
$\leq d(y_{n},Tx_{n})+d(x_{n},y_{n})$
$=d(\gamma(x_{n},Tx_{n};\alpha_{n}),Tx_{n})+d(\gamma(x_{n},Tx_{n};\alpha_{n}),x_{n})$
$=(1-\alpha_{n})d(x_{n},Tx_{n})+\alpha_{n}d(Tx_{n},Tx_{n})+(1-\alpha_{n})d(x_{n},x_{n})+\alpha_{n}d(Tx_{n},x_{n})$
$=(1-\alpha_{n})d(x_{n},Tx_{n})+\alpha_{n}d(Tx_{n},x_{n})=d(x_{n},Tx_{n}).$
(iv) Using (41), we have
$\liminf_{n\to\infty}d(x_{n},Tx_{n})=0,$ (42)
because otherwise $d(x_{n},Tx_{n})\geq K$ for all $n\in\mathbb{N}$ for some $K>0$ and then
$\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})\geq K\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty,$
which is a contradiction to the condition (41). From (iii) sequence $\{d(x_{n},Tx_{n})\}$ is decreasing. Combining this fact with (42) gives $\lim_{n\to\infty}d(x_{n},Tx_{n})=0$. Hence, $\{x_{n}\}$ satisfies the LEAF point property. Therefore, from Proposition 10, $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$. ∎
Theorem 24.
Let $C$ be a nonempty, closed and geodesically convex subset of a Hadamard manifold $\mathcal{M}$ and let $T\colon C\to C$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be the sequence generated by $p$-accelerated normal S-iteration method (38), where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ satisfying the condition $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$. Then, we have the following:
  1. (i)
    $d(x_{n},Tx_{n})\leq\frac{d(x_{0},v)}{\sqrt{\sum_{i=1}^{n}\alpha_{i}(1-\alpha_{i})}}$ and $d(x_{n},Tx_{n})=o\left(\frac{1}{\sqrt{\sum_{i=1}^{n}\alpha_{i}(1-\alpha_{i})}}\right)$.
  2. (ii)
    If in addition $0<a\leq\alpha_{n}\leq b<1$, then $d(x_{n},Tx_{n})=o\left(\frac{1}{\sqrt{n}}\right)$.
Proof.
(i) Using (40), we have
$\sum_{i=1}^{n}\alpha_{i}(1-\alpha_{i})d^{2}(x_{i},Tx_{i})\leq\sum_{i=1}^{n}(d^{2}(x_{i},v)-d^{2}(x_{i+1},v))=d^{2}(x_{0},v).$ (43)
From Theorem 23(iii), sequence $\{d(x_{n},Tx_{n})\}$ is decreasing. It follows from (43) that
$d^{2}(x_{n},Tx_{n})\sum_{i=1}^{n}\alpha_{i}(1-\alpha_{i})\leq d^{2}(x_{0},v).$
This implies that
$d(x_{n},Tx_{n})\leq\frac{d(x_{0},v)}{\sqrt{\sum_{i=1}^{n}\alpha_{i}(1-\alpha_{i})}}.$
From Theorem 23(ii), we have $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})<\infty$. Since $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$, it follows from Lemma 1 that
$d(x_{n},Tx_{n})=o\left(\frac{1}{\sqrt{\sum_{i=1}^{n}\alpha_{i}(1-\alpha_{i})}}\right).$
(ii) Since $0<a\leq\alpha_{n}\leq b<1$, from Theorem 23(ii), we have $\sum_{n=1}^{\infty}a(1-b)d^{2}(x_{n},Tx_{n})<\infty$ or $\sum_{n=1}^{\infty}d^{2}(x_{n},Tx_{n})<\infty$. Again it follows from Lemma 1 that
$d(x_{n},Tx_{n})=o\left(\frac{1}{\sqrt{n}}\right).$
Remark 25.
While both Inertial Mann method (11) and $p$-accelerated normal S-iteration method (38) aim to compute a fixed point of the nonexpansive mapping $T$, their convergence behavior differs significantly in practice. The $p$-accelerated normal S-iteration method (38) generates a relatively stable sequence, and the fixed-point residual $d(Tx_{n},x_{n})$ typically decreases in a monotonic fashion. As such, the value of $d(Tx_{n},x_{n})$ itself serves as a reliable indicator of the convergence rate.
On the other hand, the inertial Mann iteration method (11), which incorporates momentum via previous iterates, produce more oscillatory trajectories due to the presence of momentum terms. As a result, the standard residual $\{d(Tx_{n},x_{n})\}$ may not monotonically decrease, and may not reflect the actual rate at which the iterates approach the fixed point set. To better assess convergence behavior, we therefore consider the best iterate criterion $R_{T,\{x_{n}\}}(n)=\underset{{1\leq i\leq n}}{\min}d(Tx_{i},x_{i})$, which is more appropriate for methods with inertial components.
Now we introduce two accelerated Douglas-Rachford-type algorithms for solving the minimization problem (2) on Hadamard manifolds. By defining $T=R_{\lambda\phi}R_{\lambda\psi}$, the classical Douglas-Rachford method can be seen as a special case of the Krasnoselskii-Mann iteration method. The proposed algorithms extend this idea to Hadamard manifolds and incorporate acceleration techniques. Specifically, both algorithms are constructed using the composition of the reflected resolvents $R_{\lambda\phi}$ and $R_{\lambda\psi}$, following the structure of the inertial Mann iteration method (11) and the $p$-accelerated normal S-iteration method (38).
Algorithm 26.
[…]
Initialization: Choose $x_{0},x_{1}\in\mathcal{M}$ arbitrary.
Iterative step: For the current $n^{th}$ iterate, compute the $(n+1)^{th}$ iteration as follows:
$\begin{cases}y_{n}=\exp_{x_{n}}(-\theta_{n}\exp_{x_{n}}^{-1}x_{n-1}),\\ x_{n+1}=\gamma(y_{n},{R}_{\lambda\phi}{R}_{\lambda\psi}(y_{n});\alpha_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (44)
where $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ are sequences in $(0,1)$ and $\lambda>0$.
Algorithm 27.
[…]
Initialization: Choose $x_{1}\in\mathcal{M}$ arbitrary.
Iterative step: For the current $n^{th}$ iterate, compute the $(n+1)^{th}$ iteration as follows:
$\begin{cases}y_{n}=\gamma(x_{n},{R}_{\lambda\phi}{R}_{\lambda\psi}(x_{n});\alpha_{n}),\\ x_{n+1}=({R}_{\lambda\phi}{R}_{\lambda\psi})^{p}(y_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (45)
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ and $\lambda>0$.
We refer to the method defined by (44) as the inertial Douglas-Rachford method, and the method defined by (45) as the $p$-accelerated normal S-Douglas-Rachford method.
Remark 28.
If $\mathcal{M}=\mathbb{R}^{m}$, the Euclidean space, and we choose $A=\partial\phi$ and $B=\partial\psi$ in the inertial Douglas-Rachford algorithm of Boţ et al. [Bot2015, Theorem 8], then their method is equivalent to Algorithm 26. Note that we use parameter conditions different from those in Boţ et al. [Bot2015]; see Remark 17.
We now study the convergence analysis of the inertial Douglas-Rachford method (44) and the $p$-accelerated normal S-Douglas-Rachford method (45), utilizing the above results established for nonexpansive operators.
Theorem 29.
[…]
Proof.
Since $T={R}_{\lambda\phi}{R}_{\lambda\psi}$ is nonexpansive. From Theorem 18, sequence $\{x_{n}\}$ generated by (44) converges to an element $v\in\operatorname*{Fix}({R}_{\lambda\phi}{R}_{\lambda\psi})$. From Theorem 12, $u=\operatorname*{Prox}_{\lambda\psi}(v)$ is solution of problem (2). ∎
Theorem 30.
[…]
Proof.
Since $T={R}_{\lambda\phi}{R}_{\lambda\psi}$ is nonexpansive. From Theorem 23, sequence $\{x_{n}\}$ generated by (45) converges to an element $v\in\operatorname*{Fix}({R}_{\lambda\phi}{R}_{\lambda\psi})$. From Theorem 12, $u=\operatorname*{Prox}_{\lambda\psi}(v)$ is solution of problem (2). ∎
4 Numerical Experiments
In this section, we perform numerical experiments on a well-known nonconvex optimization problem, the Rosenbrock problem, defined as follows:
$\min_{x\in\mathbb{R}^{2}}a(x_{1}^{2}-x_{2})^{2}+(x_{1}-b)^{2},$ (46)
where $a,b\in(0,\infty)$ (see [Rosenbrock]). In order to handle the nonconvex problem (46) in $\mathcal{M}=\mathbb{R}^{2}$, consider
$G_{x}=\begin{pmatrix}1+4x_{1}^{2}&-2x_{1}\\ -2x_{1}&1\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathbb{R}^{2},$
which has the inverse matrix
$G_{x}^{-1}=\begin{pmatrix}1&2x_{1}\\ 2x_{1}&1+4x_{1}^{2}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathbb{R}^{2}.$
Consider the inner product on $\mathcal{T}_{x}\mathcal{M}=\mathbb{R}^{2}$ defined by
$\langle u,v\rangle_{x}=u^{T}G_{x}v\text{ for all }u,v\in\mathcal{T}_{x}\mathcal{M}.$ (47)
Then $\mathcal{M}$ is a Hadamard manifold of zero sectional curvature with respect to metric (47) (see [DaCruz]) and the following statements hold:
  1. (a)
    The Riemannian distance between any $x=(x_{1},x_{2})^{T}$ and $y=(y_{1},y_{2})^{T}$ in $\mathcal{M}$ is given by
    $d(x,y)=\left((x_{1}-y_{1})^{2}+(x_{1}^{2}-y_{1}^{2}-x_{2}+y_{2})^{2}\right)^{1/2}.$
  2. (b)
    The geodesic $\gamma(x,y;\cdot)\colon[0,1]\to\mathcal{M}$ joining the points $x=(x_{1},x_{2})^{T}$ and $y=(y_{1},y_{2})^{T}$ is
    $\gamma(x,y;t)=\begin{pmatrix}x_{1}+t(y_{1}-x_{1})\\ (x_{2}+t((y_{2}-x_{2})-(y_{1}-x_{1}^{2})+t^{2}(y_{1}-x_{1})^{2})\end{pmatrix}$.
  3. (c)
    For $x=(x_{1},x_{2})^{T}$, the exponential map $\exp_{x}\colon\mathcal{T}_{x}\mathcal{M}\to\mathcal{M}$ is given by
    $\exp_{x}(u)=\begin{pmatrix}x_{1}+u_{1}\\ x_{2}+u_{2}+u_{1}^{2}\end{pmatrix}\text{ for all }u=(u_{1},u_{2})^{T}\in\mathcal{T}_{x}\mathcal{M},$
    which has the inverse $\exp_{x}^{-1}\colon\mathcal{M}\to\mathcal{T}_{x}\mathcal{M}$ given by
    $\exp_{x}^{-1}(y)=\begin{pmatrix}y_{1}-x_{1}\\ y_{2}-x_{2}-(y_{1}-x_{1})^{2}\end{pmatrix}\text{ for all }y=(y_{1},y_{2})^{T}\in\mathcal{M}.$
Remark 31.
  1. (i)
    If $f\colon\mathcal{M}\to\mathbb{R}$ is a real-valued function, then, the Riemannian gradient of $f$, $\operatorname*{grad}f\colon\mathcal{M}\to T\mathcal{M}$, is given by
    $\operatorname*{grad}f(x)=G_{x}^{-1}\nabla f(x)\text{ for all }x\in\mathcal{M}.$
  2. (ii)
    If $\Phi\colon\mathbb{R}^{2}\to\mathcal{M}$ is an operator defined by
    $\Phi(x)=(x_{1},x_{1}^{2}-x_{2})\text{ for all }x=(x_{1},x_{2})^{T}\in\mathbb{R}^{2},$
    then $\Phi$ is an isometry between the Euclidean space $\mathbb{R}^{2}$ and $\mathcal{M}$.
For $a,b\in(0,\infty)$, consider functions $\phi,\psi\colon\mathcal{M}\to\mathbb{R}$ defined by
$\phi(x)=a(x_{1}^{2}-x_{2})^{2}\text{ and }\psi(x)=(x_{1}-b)^{2}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}.$ (48)
Using the isometry $\Phi$, we get
$\phi\circ\Phi(x)=a(x_{1}^{2}-x_{1}^{2}+x_{2})^{2}=ax_{2}^{2}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}$
and
$\psi\circ\Phi(x)=(x_{1}-b)^{2}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}.$
Thus $\phi$ and $\psi$ are geodesically convex on $\mathcal{M}$. For more details of this splitting; see [BergmannJOTA2024].
We observe that
  1. (i)
    The Euclidean gradients of $\phi$ and $\psi$ are given by
    $\nabla\phi(x)=\begin{pmatrix}4ax_{1}(x_{1}^{2}-x_{2})\\ -2a(x_{1}^{2}-x_{2})\end{pmatrix}\text{ and }\nabla\psi(x)=\begin{pmatrix}2(x_{1}-b)\\ 0\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathbb{R}^{2}.$
  2. (ii)
    In view of Remark 31(i), the gradient of $\phi$ and $\psi$ are given by
    $\operatorname*{grad}\phi(x)=\begin{pmatrix}0\\ -2a(x_{1}^{2}-x_{2})\end{pmatrix}\text{ and }\operatorname*{grad}\psi(x)=\begin{pmatrix}2(x_{1}-b)\\ 4x_{1}(x_{1}-b)\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}.$
  3. (iii)
    For $\lambda>0$, the proximal operators $\operatorname*{Prox}_{\lambda\phi}$ and $\operatorname*{Prox}_{\lambda\psi}$ are defined by
    $\operatorname{Prox}_{\lambda\phi}(x)=\begin{pmatrix}x_{1}\\ \frac{x_{2}+2a\lambda x_{1}^{2}}{1+2a\lambda}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}\text{ and }$
    $\operatorname{Prox}_{\lambda\psi}(x)=\begin{pmatrix}\frac{x_{1}+2\lambda b}{1+2\lambda}\\ x_{2}-\frac{4\lambda(x_{1}+2\lambda b)(x_{1}-b)+4\lambda^{2}(x_{1}-b)^{2}}{(1+2\lambda)^{2}}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}.$
    For $\lambda>0$, the reflection $R_{\lambda\phi}$ of $\operatorname*{Prox}_{\lambda\phi}$ and the reflection $R_{\lambda\psi}$ of $\operatorname*{Prox}_{\lambda\psi}$ are defined by
    $R_{\lambda\phi}(x)=\begin{pmatrix}x_{1}\\ \frac{2(x_{2}+2a\lambda x_{1}^{2})}{1+2a\lambda}-x_{2}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}\text{ and }$ (49)
    $R_{\lambda\psi}(x)=\begin{pmatrix}\frac{(1-2\lambda)x_{1}+4\lambda b}{1+2\lambda}\\ x_{2}-\frac{8\lambda(x_{1}+2\lambda b)(x_{1}-b)}{(1+2\lambda)^{2}}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}.$ (50)
By Proposition 15, the reflections $R_{\lambda\phi}$ and $R_{\lambda\psi}$ defined by (49) and (50), respectively, are nonexpansive.
Note that, the Rosenbrock problem (46) can be transformed into the following minimization problem:
$\min_{x\in\mathcal{M}}\phi(x)+\psi(x),$ (51)
where $\phi$ and $\psi$ are defined by (48).
For the numerical experiment, we compare three different methods: the Douglas-Rachford method [Alg(DR)] (3), the inertial Douglas-Rachford method [Alg(InDR)] (44) and the $p$-accelerated normal S-Douglas-Rachford method [Alg($p$-AccDR)] (45). All numerical tests are carried out using MATLAB R2024b on an Apple Macbook Air Laptop with a 1.8 GHz Dual-Core Intel Core i5 processor and 8 GB RAM. All algorithms stops when they meet the stopping criterion $E_{n}=d(s_{n},s_{n-1})<10^{-14}$, where $s_{n}=\operatorname*{Prox}_{\lambda\psi}(x_{n})$.
Numerical experiments are performed on the Rosenbrock function (46) with parameters $(a,b)=(1,2)$ and $(3,7)$. The initial point is chosen as $x_{0}=(1,2)^{\top}$ for all methods, while Method (InDR) uses $x_{1}=(1,3)^{\top}$ in both cases.
We set $\lambda=1$ for all three algorithms and iteration parameters are chosen as follows:
Algorithm $(a,b)=(1,2)$ $(a,b)=(1,2)$ $(a,b)=(3,7)$ $(a,b)=(3,7)$
Alg(DR) $\alpha_{n}=0.1$ $\alpha_{n}=0.3$ $\alpha_{n}=0.2$ $\alpha_{n}=0.4$
Alg(InDR) $\alpha_{n}=0.1,$ $\alpha_{n}=0.3,$ $\alpha_{n}=0.2,$ $\alpha_{n}=0.4,$
$\theta_{n}=0.42$ $\theta_{n}=0.35$ $\theta_{n}=0.4$ $\theta_{n}=0.33$
Alg($p$-AccDR) $\alpha_{n}=0.1$ $\alpha_{n}=0.3$ $\alpha_{n}=0.2$ $\alpha_{n}=0.4$
Note that in our parameter choice, we have taken the same $\alpha_{n}$ (in the Mann-type step) for all methods. Corresponding to this choice of $\alpha_{n}$, we selected the parameter $\theta_{n}$ according to condition (C3) for Alg(InDR), that is, $\epsilon=(1-\alpha_{n})/\alpha_{n}$, $\epsilon_{1}=1+\epsilon+\max\{1,\epsilon\}$, and $\theta_{n}\approx 0.9(\epsilon/\epsilon_{1})$. We chose $p=2$, which is the smallest value that yielded the best performance in our experiments. These parameters were chosen empirically by testing several reasonable values.
The runtime and number of iterations are presented in Table 1 for $(a,b)=(1,2)$ and in Table 2 for $(a,b)=(3,7)$. The graph of $E_{n}$ values with respect to the number of iterations $n$ is shown in Figure 2 for both cases. From Tables 1, 2, and Figure 2, it is evident that Alg($p$-AccDR) is faster than Alg(InDR) and Alg(DR) in terms of both runtime and number of iterations.
Table 1: Numerical performance of the runtime and number of iterations for Rosenbrock problem with $(a,b)=(1,2)$
$(\alpha_{n},\theta_{n})=(0.1,0.42)$ $(\alpha_{n},\theta_{n})=(0.3,0.35)$
Algorithm Runtime(sec) Iterations Runtime(sec) Iterations
Alg(DR) 0.000800 213 0.000468 63
Alg(InDR) 0.000461 102 0.000287 43
Alg(pp-AccDR) 0.000088 15 0.000099 13
Table 2: Numerical performance of the runtime and number of iterations for Rosenbrock problem with $(a,b)=(3,7)$
$(\alpha_{n},\theta_{n})=(0.2,0.4)$ $(\alpha_{n},\theta_{n})=(0.4,0.33)$
Algorithm Runtime(sec) Iterations Runtime(sec) Iterations
Alg(DR) 0.001003 106 0.000623 44
Alg(InDR) 0.000385 55 0.000368 37
Alg(pp-AccDR) 0.000150 30 0.000130 19
(a) $(a,b)=(1,2),(\alpha_{n},\theta_{n})=(0.1,0.42)$
(b) $(a,b)=(1,2),(\alpha_{n},\theta_{n})=(0.3,0.35)$
(c) $(a,b)=(3,7),(\alpha_{n},\theta_{n})=(0.2,0.4)$
(d) $(a,b)=(3,7),(\alpha_{n},\theta_{n})=(0.4,0.33)$
Figure 2: Comparison of $E_{n}$ with respect to number of iterations for the Rosenbrock Problem
5 Applications
In this section, we present applications of the proposed algorithms for solving the problem of minimizing a functional involving multiple summands. To this end, we develop parallel versions of the proposed algorithms and apply them to find the solution of generalized Heron problem on Hadamard manifolds.
5.1 Minimization of functional containing multiple summands
[…]
$\min_{x\in\mathcal{M}}\sum_{k\in\varLambda_{N}}f_{k}(x).$ (52)
Now, consider the product manifold $\mathcal{M}^{N}=\{\mathbf{x}=(x_{1},\ldots,x_{N})\colon x_{k}\in\mathcal{M},k\in\varLambda_{N}\}$. Define
$\mathbf{D}=\{\mathbf{x}\in\mathcal{M}^{N}\colon x_{1}=\dots=x_{N}\in\mathcal{M}\}.$ (53)
Then $\mathbf{D}$ is a nonempty, closed and geodesically convex set. Hence, $\iota_{\mathbf{D}}$ is a proper, geodesically convex and lower semicontinuous function; see [BacakNAA2014]. For $\mathbf{x}=(x_{1},\ldots,x_{N})\in\mathcal{M}^{N}$
$\operatorname{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{x})=P_{\mathbf{D}}(\mathbf{x})=\Bigl(\operatorname*{argmin}_{x\in\mathcal{M}}\sum_{k\in\varLambda_{N}}d^{2}(x_{k},x),\ldots,\operatorname*{argmin}_{x\in\mathcal{M}}\sum_{k\in\varLambda_{N}}d^{2}(x_{k},x)\Bigl)\in\mathcal{M}^{N}.$
Define $F\colon\mathcal{M}^{N}\to(-\infty,\infty]$ by
[…]
We are interested in minimization of $F$ over $\mathbf{D}$. For this, we consider the following unconstrained minimization problem:
$\min_{\mathbf{x}\in\mathcal{M}^{N}}F(\mathbf{x})+\iota_{\mathbf{D}}(\mathbf{x}).$ (55)
A detailed explanation of the above formulation is given in [Bergmann2016]. Note that for $\lambda>0$
$\operatorname{Prox}_{\lambda F}(\mathbf{x})$ $=\left(\operatorname*{Prox}_{\lambda f_{1}}(x_{1}),\ldots,\operatorname*{Prox}_{\lambda f_{N}}(x_{N})\right)\text{ for all }\mathbf{x}\in\mathcal{M}^{N}.$
It follows that the reflection of $F$ is just given by the componentwise reflections of $\operatorname*{Prox}_{\lambda f_{k}}$, i.e.,
$R_{\lambda F}(\mathbf{x})=\left(R_{\lambda f_{1}}(x_{1}),\ldots,R_{\lambda f_{N}}(x_{N})\right)\text{ for all }\mathbf{x}\in\mathcal{M}^{N}.$ (56)
We now present two methods based on the inertial Douglas-Rachford method (44) and the $p$-accelerated normal S-Douglas-Rachford method (45) for finding a solution to problem (52).
Corollary 32.
[…]
$\begin{cases}\mathbf{y}_{n}=\exp_{\mathbf{x}_{n}}(-\theta_{n}\exp_{\mathbf{x}_{n}}^{-1}\mathbf{x}_{n-1}),\\ \mathbf{x}_{n+1}=\gamma(\mathbf{y}_{n},{R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}}(\mathbf{y}_{n});\alpha_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (57)
where $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ are sequences satisfying the conditions (C1), (C2) and (C3). Then $\{\mathbf{x}_{n}\}$ converges to an element $\mathbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}})$ such that $\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{v})$ is solution of problem (55).
Proof.
Since $T=R_{\lambda F}R_{\lambda\iota_{\mathbf{D}}}$ is nonexpansive. From Theorem 29, sequence $\{\mathbf{x}_{n}\}$ generated by (57) converges to an element $\mathbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}})$. From Theorem 12, we have $\mathbf{u}=\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{v})$ is solution of problem (55). ∎
Corollary 33.
[…]
$\begin{cases}\mathbf{y}_{n}=\gamma(\mathbf{x}_{n},{R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}}(\mathbf{x}_{n});\alpha_{n}),\\ \mathbf{x}_{n+1}=({R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}})^{p}(\mathbf{y}_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (58)
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ satisfying the condition $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$. Then $\{\mathbf{x}_{n}\}$ converges to an element $\mathbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}})$ such that $\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{v})$ is solution of problem (55).
Proof.
Since $T=R_{\lambda F}R_{\lambda\iota_{\mathbf{D}}}$ is nonexpansive. From Theorem 30, sequence $\{\mathbf{x}_{n}\}$ generated by (58) converges to an element $\mathbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}})$. From Theorem 12, we have $\mathbf{u}=\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{v})$ is solution of problem (55). ∎
We refer to the method defined by (57) as the inertial parallel Douglas-Rachford method, and the method defined by (58) as the $p$-accelerated normal S-parallel Douglas-Rachford method.
5.2 The generalized Heron problem in Hadamard manifolds
Heron from Alexandria (10-75 AD) was “a Greek geometer and inventor whose writings preserved for posterity a knowledge of the mathematics and engineering of Babylonia, ancient Egypt, and the Greco-Roman world” (from the Encyclopedia Britannica). One of the geometric problems he proposed in his Catroptica was as follows: find a point on a straight line in the plane such that the sum of the distances from it to two given points is minimal.
In 2012, Mordukhovich et al. [Mordukhovich2012] generalized the Heron problem. In the examination of the generalized Heron problem, as discussed in reference [Mordukhovich2012, Mordukhovich2012(2), Bot2015, Dixit2022, ShikherOpt2024, ReichFT2023], the conventional notions of points and straight lines are substituted with nonempty, closed and convex sets. Within this context, the objective is to locate a point within a predetermined nonempty, closed and convex set $C\subset\mathbb{R}^{m}$ in such a way that the combined distance to specified nonempty, closed and convex sets $C_{k}\subset\mathbb{R}^{m},k\in\varLambda_{N}$ is minimized. Mathematically, we can write
$\text{minimize }\sum_{k\in\varLambda_{N}}d(x,C_{k})\text{ subject to }x\in C\subset\mathbb{R}^{m}.$ (59)
Now, we consider the generalized Heron problem in the setting of Hadamard manifold $\mathcal{M}$, where the objective is to locate a point within a predetermined nonempty, closed and geodesically convex set $C\subset\mathcal{M}$ in such a way that the combined distance to specified nonempty, closed and geodesically convex sets $C_{k}\subset\mathcal{M},k\in\varLambda_{N}$ is minimized. Mathematically, the generalized Heron problem can be formulated as:
$\text{minimize }\sum_{k\in\varLambda_{N}}d(x,C_{k})\text{ subject to }x\in C\subset\mathcal{M}.$ (60)
The constrained problem (59) can be formulated into the following unconstrained problem:
$\min_{x\in\mathcal{M}}\sum_{k\in\varLambda_{N}}\phi_{k}(x)+\iota_{C}(x),$ (61)
where $\phi_{k}(x)=d(x,C_{k})$ for all $k\in\varLambda_{N}$. In the view of (55), problem (61) can be written as
$\min_{\mathbf{x}\in\mathcal{M}^{N+1}}F(\mathbf{x})+\iota_{\mathbf{D}}(\mathbf{x}),$ (62)
where $F(\mathbf{x})=\sum_{k\in\varLambda_{N}}d(x_{k},C_{k})+\iota_{C}(x_{N+1})$ and $\mathbf{D}$$\{\mathbf{x}\in\mathcal{M}^{N+1}\colon x_{1}=\dots=x_{N+1}\}$.
Remark 34.
By (56), the reflection of $F$ defined in (62) is given by
$R_{\lambda F}(\mathbf{x})=\bigl(R_{\lambda d(x_{1},C_{1})}(x_{1}),\ldots,R_{\lambda d(x_{N},C_{N})}(x_{N}),R_{\lambda\iota_{C}}(x_{N+1})\bigr)\quad\text{for all }\mathbf{x}\in\mathcal{M}^{N+1}.$
Since $\mathcal{M}^{N+1}$ has constant sectional curvature, it follows from Proposition 13 and Theorem 14 that the reflection $R_{\lambda F}$ is nonexpansive. Similarly, by Proposition 13, the reflection $R_{\lambda\iota_{\mathbf{D}}}$ is nonexpansive, since $\mathbf{D}$ is a nonempty, closed, and geodesically convex set.
Example 35.
Let $\mathcal{M}=\mathbb{R}^{m}_{++}=\{x=(x_{1},\ldots,x_{m})\in\mathbb{R}^{m}\colon x_{i}>0,i=1,\ldots,m\}$. For $x\in\mathcal{M}$, let
$G(x)=\operatorname{diag}(1/x_{1}^{2},\ldots,1/x_{m}^{2})\text{ and }H(x)=\operatorname{diag}(x_{1},\ldots,x_{m})$
be two diagonal matrices. Then $\mathcal{M}$ is a Riemannian manifold with the Riemannian metric defined by
$\langle u,v\rangle_{x}=\langle u,G(x)v\rangle\text{ for }x\in\mathcal{M}\text{ and }u,v\in\mathcal{T}_{x}\mathcal{M}=\mathbb{R}^{m}.$ (63)
The Riemannian distance $d\colon\mathcal{M}\times\mathcal{M}\to[0,\infty)$ is given by
$d(x,y)=\left(\sum_{i=1}^{m}\left(\ln\left(\frac{x_{i}}{y_{i}}\right)\right)^{2}\right)^{1/2}\text{ for all }x=(x_{1},\ldots,x_{m}),y=(y_{1},\ldots,y_{m})\in\mathcal{M}.$
$\mathcal{M}$ is a Hadamard manifold with zero sectional curvature (see [DaCruz]). For $x=(x_{1},\ldots,x_{m}),y=(y_{1},\ldots,y_{m})\in\mathcal{M}$, the geodesic joining $x$ to $y$ is given by
$\gamma(x,y;t)=(x_{1}^{1-t}y_{1}^{t},\ldots,x_{m}^{1-t}y_{m}^{t})\text{ for all }t\in[0,1];$
the exponential map $\exp_{x}\colon\mathcal{T}_{x}\mathcal{M}\to\mathcal{M}$ is given by
$\exp_{x}(v)=\left(x_{1}e^{(v_{1}/x_{1})},\ldots,x_{m}e^{(v_{m}/x_{m})}\right)\text{ for all }v=(v_{1},\ldots,v_{m})\in\mathcal{T}_{x}\mathcal{M};$
the inverse of the exponential map $\exp_{x}^{-1}\colon\mathcal{M}\to\mathcal{T}_{x}\mathcal{M}$ is given by
$\exp^{-1}_{x}y=\left(x_{1}\ln\left({y_{1}}/{x_{1}}\right),\ldots,x_{m}\ln\left({y_{m}}/{x_{m}}\right)\right)\text{ for all }y=(y_{1},\ldots,y_{m})\in\mathcal{M}.$
We denote the closed ball centered at $c\in\mathcal{M}$ with radius $r$ by $B_{r}[c]$, i.e., $B_{r}[c]=\{x\in\mathcal{M}:d(x,c)\leq r\}.$
Remark 36.
For the Hadamard manifold $\mathcal{M}=\mathbb{R}_{++}^{m}$ with the Riemannian metric defined by (63), we have the following:
  1. (a)
    $\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{x})=P_{\mathbf{D}}(\mathbf{x})=\left(b,\ldots,b\right)\in\mathcal{M}^{N}$ for all $\mathbf{x}=(x_{1},\ldots,x_{N})\in\mathcal{M}^{N}$, where $x_{k}=(x_{k,1},x_{k,2},\ldots,x_{k,m})$, $k\in\varLambda_{N}$ and $b=\left(\left(\prod_{k\in\varLambda_{N}}x_{k,1}\right)^{\frac{1}{N}},\ldots,\left(\prod_{k\in\varLambda_{N}}x_{k,m}\right)^{\frac{1}{N}}\right)$.
  2. (b)
    $P_{B_{r}[c]}(x)=\begin{cases}\gamma\left(c,x;\frac{r}{d(c,x)}\right),&\text{ if }d(c,x)>r,\\ x,&\text{ if }d(c,x)\leq r.\end{cases}$
  3. (c)
    $\operatorname*{Prox}_{\lambda d(x,c)}(x)=\begin{cases}\gamma(x,c;\theta),&\text{ if }d(x,c)>\lambda,\\ c,&\text{ if }d(x,c)\leq\lambda,\end{cases}$ where $\theta=\frac{\lambda}{d(x,c)}$.
  4. (d)
    $\operatorname*{Prox}_{\lambda d(x,B_{r}[c])}(x)=\begin{cases}\gamma(x,P_{B_{r}[c]}(x);\theta),&\text{ if }d(x,B_{r}[c])>\lambda,\\ P_{B_{r}[c]}(x),&\text{ if }d(x,B_{r}[c])\leq\lambda,\end{cases}$ where $\theta=\frac{\lambda}{d(x,B_{r}[c])}$.
For the numerical experiment, we first consider the generalized Heron problem for a ball with point constraints in $\mathcal{M}=\mathbb{R}_{++}^{m}$ with the Riemannian metric defined by (63), i.e.,
$\text{minimize }\sum_{k\in\varLambda_{N}}d(x,a_{k})\text{ subject to }x\in C\subset\mathcal{M},$ (64)
where $a_{k}\in\mathcal{M},k\in\varLambda_{N}$.
We compute the solution of the generalized Heron problem using parallel Douglas-Rachford method [Bergmann2016, Algorithm 4], inertial parallel Douglas-Rachford method (58) and $p$-accelerated normal S-parallel Douglas-Rachford method (57) and compare based on their efficiency.
By using Remark 36(a), we may assume that $\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{x}_{n})=(t_{n},\ldots,t_{n})\in\mathcal{M}^{N+1}$. We then define the error term by $Er(n)=d(t_{n},t_{n-1})$. In all experiments, the algorithm is terminated when the stopping criterion $Er(n)\leq 10^{-10}$ is satisfied.
We take $\lambda=1$ and iteration parameters for these algorithms are listed below:
[Bergmann2016, Algorithm 4][Alg(PDRA)] $:$ $\alpha_{n}=0.4;$
Algorithm (58)[Alg(In-M)] $:$ $\alpha_{n}=0.4,\theta_{n}=0.08;$
Algorithm (57)[Alg($p$-Acc)] $:$ $\alpha_{n}=0.4.$
Note that we use the same parameter $\alpha_{n}$ for all the algorithms. However, the inertial parameter $\theta_{n}$ is chosen according to condition (C3), that is, $\epsilon=(1-\alpha_{n})/\alpha_{n}$, $\epsilon_{1}=1+\epsilon+\max\{1,\epsilon\}$, and $\theta_{n}\approx 0.25(\epsilon/\epsilon_{1})$. We chose $p=2$, which is the smallest value that yielded the best performance in our experiments. The parameters were chosen by testing several reasonable values, and those giving the best convergence are reported.
Example 37.
Consider the generalized Heron problem (64), where $m=2,N=2$, the constraint set $C=B_{r}[c]$ is a ball in $\mathcal{M}$ centered at $c=(35,35)$ with radius $r=0.4$. Consider two cases:
Case 1: Let $a_{1}=(15,70)$ and $a_{2}=(70,15)$.
Case 2: Let $a_{1}=(5,50)$ and $a_{2}=(50,5)$.
Note that in both cases, $C$ does not contain $a_{1}$ and $a_{2}$. Now we have two observations:
  1. (i)
    In Case 1, the geodesic segment $\gamma(a_{1},a_{2};\cdot)$ intersects the ball $C$; see Figure 3(a). Then any point in the intersection is a solution of problem (64). In this case, problem (64) has infinitely many solutions.
  2. (ii)
    In Case 2, the geodesic segment $\gamma(a_{1},a_{2};\cdot)$ does not intersect the ball $C$; see Figure 3(b). Then, there is a unique point on the boundary of the ball $C$.
(a) Intersects
(b) Does not intersect
Figure 3: Generalized Heron problems for two points with a ball constraint.
We take the initial point $\mathbf{x}_{0}=ones(1,6)$ for all the algorithms and $\mathbf{x}_{1}=2*ones(1,6)$ for Alg(In-M) for both cases (Case 1 and Case 2) in Example 37. The numerical performance of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) are depicted in Table 3. We plot the relationship between the number of iterations and $Er(n)$; see Figure 4. From Table 3 and Figure 4, we observe that Alg($p$-Acc) requires less computational time and fewer iterations to achieve the desired stopping criterion in both cases.
Table 3: Numerical performance in terms of CPU time and number of iterations for Alg(PDRA), Alg(In-M), and Alg($p$-Acc) for Example 37
Case 1 Case 2
Algorithm CPU(s) iter(n) $Er(n)$ CPU(s) iter(n) $Er(n)$
Alg(PDRA) $0.029219$ $75$ $7.4085*10^{-11}$ $0.029682$ $99$ $9.7691*10^{-11}$
Alg(In-M) $0.015613$ $46$ $9.7055*10^{-11}$ $0.022472$ $66$ $7.3605*10^{-11}$
Alg(pp-Acc) $0.012893$ 1919 $6.3342*10^{-11}$ $0.012026$ $31$ $3.6084*10^{-11}$
(a) Case 1
(b) Case 2
Figure 4: Comparison of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) based on the relationship between the number of iterations and $Er(n)$ for Example 37.
Example 38.
Consider the generalized Heron problem (64) for the following two cases:
Case 1: $m=2$, $N=4$.
We take predetermined set $C$ as ball $B_{r}[c]$ with $c=(35,35)$ and $r=0.4$ and choose constraints sets to be points $a_{1}=(15,15)$, $a_{2}=(65,65)$, $a_{3}=(10,60)$ and $a_{4}=(60,10)$; see Figure 5.
Case 2: $m=20$, $N=10$.
We take predetermined set $C$ as ball $B_{r}[c]$ with
$c=(41,29,61,39,121,79,99,51,81,59,111,39,79,31,61,19,121,61,81,39)\text{ and }r=0.4$
and choose constraints sets to be points $a_{k}$, $k=1,2,\ldots,10$, where
$a_{1}$ $=(35,30,60,40,120,80,100,50,80,60,110,40,80,30,60,20,120,60,80,40),$
$a_{2}$ $=(45,30,60,40,120,80,100,50,80,60,110,40,80,30,60,20,120,60,80,40),$
$a_{3}$ $=(35,30,55,40,120,80,100,50,80,60,110,40,80,30,60,20,120,60,80,40),$
$a_{4}$ $=(35,30,65,40,120,80,100,50,80,60,110,40,80,30,60,20,120,60,80,40),$
$a_{5}$ $=(35,30,60,40,115,80,100,50,80,60,110,40,80,30,60,20,120,60,80,40),$
$a_{6}$ $=(35,30,65,40,125,80,100,50,80,60,110,40,80,30,60,20,120,60,80,40),$
$a_{7}$ $=(35,30,65,40,120,80,95,50,80,60,110,40,80,30,60,20,120,60,80,40),$
$a_{8}$ $=(35,30,65,40,120,80,105,50,80,60,110,40,80,30,60,20,120,60,80,40),$
$a_{9}$ $=(35,30,65,40,120,80,100,50,75,60,110,40,80,30,60,20,120,60,80,40),$
$a_{10}$ $=(35,30,65,40,120,80,100,50,85,60,110,40,80,30,60,20,120,60,80,40).$
Figure 5: The generalized Heron problem for four points with a ball constraint
We take the initial point $\mathbf{x}_{0}=ones(1,m*(N+1))$ for all the algorithms and $\mathbf{x}_{1}=2*ones(1,m*(N+1))$ for Alg(In-M) for both cases (Case 1 and Case 2) in Example 38. The numerical performance of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) are depicted in Table 4 for both cases (Case 1 and Case 2). We plot the relationship between the number of iterations and $Er(n)$ to have a clear visualization of the experiment; see Figure 6. From Table 4 and Figure 6, we see that our Algorithm Alg($p$-Acc) requires less computational time and less number of iterations to achieve the desired stopping criterion in both cases. From Figure 7, we can see the convergence behaviour of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) to achieving the desired goal as number of iteration increases upto $30$th iteration for Case 2.
Table 4: Numerical performance in terms of CPU time and number of iterations for Alg(PDRA), Alg(In-M), and Alg($p$-Acc) for Example 38
Case 1 Case 2
Algorithm CPU(s) iter(n) $Er(n)$ CPU(s) iter(n) $Er(n)$
Alg(PDRA) $0.029219$ $113$ $5.2545*10^{-11}$ $0.029682$ $192$ $9.3761*10^{-11}$
Alg(In-M) $0.015613$ $83$ $4.4691*10^{-11}$ $0.022472$ $135$ $9.0710*10^{-11}$
Alg(pp-Acc) $0.012893$ $34$ $6.0025*10^{-12}$ $0.012026$ $62$ $4.3786*10^{-11}$
(a) Case 1
(b) Case 2
Figure 6: Comparison of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) based on the relationship between the number of iterations and $Er(n)$ for Example 38.
(a) Alg(PDRA)
(b) Alg(In-M)
(c) Alg($p$-Acc)
Figure 7: Graphical progression of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) upto 30 iterations for Case 2 in Example 38.
Example 39.
Consider the generalized Heron problem (60) for the following two cases: In the first case we consider balls with ball constraint in $\mathcal{M}=\mathbb{R}_{++}^{2}$ and in the second case balls with ball constraint in $\mathcal{M}=\mathbb{R}_{++}^{3}$ as follows:
Case 1: $m=2$, $N=4$.
We take predetermined set $C$ as ball $B_{r}[c]$ with $c=(35,35)$ and $r=0.4$ and choose constraints sets to be balls $C_{k}=B_{r}[a_{k}]$, $k=1,2,3,4$ with the center $a_{1}=(15,15)$, $a_{2}=(65,65)$, $a_{3}=(10,60)$ and $a_{4}=(60,10)$ and $r=0.4$; see Figure 8(a).
Case 2: $m=3$, $N=4$.
We take set $C$ as ball $B_{r}[c]$ with $c=(35,35,35)$ and $r=0.4$ and choose constraints sets to be balls $C_{k}=B_{r}[a_{k}]$, $k=1,2,3,4$ with the center $a_{1}=(15,15,15)$, $a_{2}=(65,65,65)$, $a_{3}=(10,60,10)$ and $a_{4}=(60,10,60)$ and $r=0.4$; see Figure 8(b).
(a) Case 1
(b) Case 2
Figure 8: Generalized Heron problems for a ball with ball constraints.
We take the initial point $\mathbf{x}_{0}=ones(1,m*(N+1))$ for all the algorithms and $\mathbf{x}_{1}=2*ones(1,m*(N+1))$ for Alg(In-M) for both cases (Case 1 and Case 2) in Example 39. The numerical performance of algorithms Alg(PDRA), Alg(In-M) and Alg($p$-Acc) are depicted in Table 5 for both cases (Case 1 and Case 2). We plot the relationship between the number of iterations and $Er(n)$ to have a clear visualization of the experiment; see Figure 9. From Table 5 and Figure 9, we see that our Algorithm Alg($p$-Acc) takes less runtime and less number of iterations to achieve the desired stopping criterion in both cases.
Table 5: Numerical performance in terms of CPU time and number of iterations for Alg(PDRA), Alg(In-M), and Alg($p$-Acc) for Example 39
Case 1 Case 2
Algorithm CPU(s) iter(n) $Er(n)$ CPU(s) iter(n) $Er(n)$
Alg(PDRA) $0.0190$ $113$ $8.9175*10^{-11}$ $0.0214$ $137$ $9.6941*10^{-11}$
Alg(In-M) $0.0146$ $83$ $4.6019*10^{-11}$ $0.0105$ $100$ $7.5025*10^{-11}$
Alg(pp-Acc) $0.0074$ $34$ $8.5375*10^{-11}$ $0.0087$ $46$ $8.6438*10^{-11}$
(a) Case 1
(b) Case 2
Figure 9: Comparison of Alg(PDRA), Alg(In-M), and Alg($p$-Acc) based on the relationship between the number of iterations and $Er(n)$ for Example 39.
6 Conclusion
In this paper, we have proposed new accelerated algorithms–the inertial Mann method and the $p$-accelerated normal S-method for finding fixed points of nonexpansive operators on Hadamard manifolds. These methods were further utilized to minimize the sum of two geodesically convex functions on Hadamard manifolds. For the minimization problem, we have derived the inertial Douglas-Rachford method and the $p$-accelerated normal S-Douglas-Rachford method, which aim to improve the convergence speed of the classical Douglas-Rachford method in the context of Hadamard manifolds. We have studied the convergence analysis of these methods and analyzed their asymptotic behavior under mild conditions on the iteration parameters.
Additionally, we introduced parallel versions of these algorithms–the inertial parallel Douglas-Rachford method and the $p$-accelerated normal S-parallel Douglas-Rachford method for finding the solution of minimization problems involving sum of more than two functions. We applied these parallel methods to find the solution of the generalized Heron problem in the context of Hadamard manifolds. We conducted numerical experiments which showed the efficiency of the proposed methods.
[…]
Acknowledgements
We thank Ronny Bergmann for his insightful discussions on the parallel Douglas-Rachford methods in the context of Hadamard manifolds, and for his suggestions regarding the numerical experiments. The authors are grateful to the anonymous referees for careful reading of the manuscript and for various suggestions that helped to improve the paper.
Declarations
  • Funding: Not applicable.
  • Conflict of interest/Competing interests: The authors declare that they have no competing interests.
  • Ethics approval and consent to participate: Not applicable.
  • Consent for publication: Not applicable.
  • Data availability: Not applicable.
  • Materials availability: Not applicable.
  • Code availability: The code used for the experiments is available from the corresponding author upon reasonable request.
  • Author contribution: All authors read and approved the final manuscript.
Appendix A Proof of Proposition 9
First, we recall some basic results on Hadamard manifolds.
Lemma 40 ([Martin1]).
Let $\mathcal{M}$ be a Hadamard manifold and let $x,y,z\in\mathcal{M}$. Then, we have the following:
  1. (i)
    Let $\varsigma_{y}$ be the angle at the vertex $y$. Then $\langle\exp_{y}^{-1}x,\exp_{y}^{-1}z\rangle=d(x,y)d(y,z)\cos\varsigma_{y}$.
  2. (ii)
    $d^{2}(x,y)+d^{2}(y,z)-2\langle\exp_{y}^{-1}x,\exp_{y}^{-1}z\rangle\leq d^{2}(z,x)$.
Lemma 41 ([Martin1]).
Let $\triangle(x,y,z)$ be a geodesic triangle in a Hadamard manifold $\mathcal{M}$ and $\triangle(x^{\prime},y^{\prime},z^{\prime})$ be its comparison triangle.
  1. (a)
    Let $\varsigma_{x},\varsigma_{y},\varsigma_{z}$ (resp. $\varsigma_{x}^{\prime},\varsigma_{y}^{\prime},\varsigma_{z}^{\prime}$) be the angles of $\triangle(x,y,z)$ (resp. $\triangle(x^{\prime},y^{\prime},z^{\prime})$) at the vertices $x,y,z$ (resp. $x^{\prime},y^{\prime},z^{\prime}$). Then $\varsigma_{x}^{\prime}\geq\varsigma_{x}$ , $\varsigma_{y}^{\prime}\geq\varsigma_{y}$ and $\varsigma_{z}^{\prime}\geq\varsigma_{z}$.
  2. (b)
    Let $w$ be a point in the geodesic joining $x$ to $y$ and $w^{\prime}$ its comparison point in the interval $x^{\prime},y^{\prime}]$. Suppose that $d(w,x)=\|w^{\prime}-x^{\prime}\|$ and $d(w,y)=\|w^{\prime}-y^{\prime}\|$. Then $d(w,z)\ \leq\|w^{\prime}-z^{\prime}\|.$
Lemma 42 ( [KhammaAOT2022]).
Let $\mathcal{M}$ be an Hadamard manifold and $x,y,z\in\mathcal{M}$. Then
$\|\exp_{x}^{-1}z-P_{x,y}\exp_{y}^{-1}z\|\leq d(x,y).$
Proof of Proposition 9. Let $u=\exp_{x}(-\theta\exp_{x}^{-1}y)$. Consider the geodesic triangle $\triangle(u,x,z)$. Then, there exists the corresponding comparison triangle $\triangle(\bar{u},\bar{x},\bar{z})\subset\mathbb{R}^{2}$ such that
$d(u,x)=\|\bar{u}-\bar{x}\|,d(u,z)=\|\bar{u}-\bar{z}\|\text{ and }d(x,z)=\|\bar{x}-\bar{z}\|.$
Observe that
$d^{2}(u,z)$ $=\|\bar{u}-\bar{z}\|^{2}=\|\bar{u}-\bar{x}+\bar{x}-\bar{z}\|^{2}=\|\bar{u}-\bar{x}\|^{2}+\|\bar{x}-\bar{z}\|^{2}+2\langle\bar{u}-\bar{x},\bar{x}-\bar{z}\rangle$
$=\|\bar{u}-\bar{x}\|^{2}+\|\bar{x}-\bar{z}\|^{2}+2\langle\bar{u}-\bar{x},\bar{x}-\bar{z}\rangle+2\|\bar{x}-\bar{z}\|^{2}-2\|\bar{x}-\bar{z}\|^{2}$
$=d^{2}(u,x)+d^{2}(x,z)+2\langle\bar{u}-\bar{z},\bar{x}-\bar{z}\rangle-2d^{2}(x,z).$ (65)
Suppose that the angles at the vertices $z$ and $\bar{z}$ are denoted by $\varsigma$ and $\varsigma^{\prime}$ respectively. Then, by Lemma 41, we have $\gamma\leq\gamma^{\prime}$. From Lemma 40(i), we get
$\langle\bar{u}-\bar{z},\bar{x}-\bar{z}\rangle$ $=\|\bar{u}-\bar{z}\|\|\bar{x}-\bar{z}\|\cos(\varsigma^{\prime})=d(u,z)d(x,z)\cos(\varsigma^{\prime})$
$\leq d(u,z)d(x,z)\cos(\varsigma)=\langle\exp_{z}^{-1}u,\exp_{z}^{-1}x\rangle.$ (66)
From (65) and (66), we have
$d^{2}(u,z)\leq d^{2}(u,x)+d^{2}(x,z)+2\langle\exp_{z}^{-1}u,\exp_{z}^{-1}x\rangle-2d^{2}(x,z).$ (67)
Note that $\exp_{x}^{-1}u=-\theta\exp_{x}^{-1}y$. Hence
$d(x,u)=\|\exp_{x}^{-1}u\|=\|\exp_{x}^{-1}(\exp_{x}(-\theta\exp_{x}^{-1}y))\|=\theta\|\exp_{x}^{-1}y\|=\theta d(x,y),$
where $\theta\in[0,1)$. It follows from (67) that
$d^{2}(u,z)$ $\leq\theta^{2}d^{2}(x,y)+d^{2}(x,z)+2\langle\exp_{z}^{-1}u,\exp_{z}^{-1}x\rangle-2d^{2}(x,z).$ (68)
Taking into consideration Remark 6 and Lemma 42, from (68) we obtain
$d^{2}(u,z)$ $\leq\theta^{2}d^{2}(x,y)+d^{2}(x,z)-2d^{2}(x,z)$
$\quad+2\langle\exp_{z}^{-1}u-P_{z,x}\exp_{x}^{-1}u+P_{z,x}\exp_{x}^{-1}u,\exp_{z}^{-1}x\rangle$
$=\theta^{2}d^{2}(x,y)+d^{2}(x,z)-2d^{2}(x,z)$
$\quad+2\langle\exp_{z}^{-1}u-P_{z,x}\exp_{x}^{-1}u,\exp_{z}^{-1}x\rangle+2\langle P_{z,x}\exp_{x}^{-1}u,\exp_{z}^{-1}x\rangle$
$\leq\theta^{2}d^{2}(x,y)+d^{2}(x,z)-2d^{2}(x,z)$
$\quad+2\|\exp_{z}^{-1}u-P_{z,x}\exp_{x}^{-1}u\|\|\exp_{z}^{-1}x\|-2\langle\exp_{x}^{-1}u,\exp_{x}^{-1}z\rangle$
$\leq\theta^{2}d^{2}(x,y)+d^{2}(x,z)-2d^{2}(x,z)+2d^{2}(z,x)-2\langle\exp_{x}^{-1}u,\exp_{x}^{-1}z\rangle$
$=\theta^{2}d^{2}(x,y)+d^{2}(x,z)-2\langle\exp_{x}^{-1}u,\exp_{x}^{-1}z\rangle.$ (69)
From the fact that $\exp_{x}^{-1}u=-\theta\exp_{x}^{-1}y$, Remark 6, Lemma 42 and (69), we have
$d^{2}(u,z)$
$\leq\theta^{2}d^{2}(x,y)+d^{2}(x,z)+2\theta\langle\exp_{x}^{-1}y,\exp_{x}^{-1}z\rangle$
$=\theta^{2}d^{2}(x,y)+d^{2}(x,z)+2\theta\langle\exp_{x}^{-1}y-P_{x,z}\exp_{z}^{-1}y+P_{x,z}\exp_{z}^{-1}y,\exp_{x}^{-1}z\rangle$
$=\theta^{2}d^{2}(x,y)+d^{2}(x,z)+2\theta\langle\exp_{x}^{-1}y-P_{x,z}\exp_{z}^{-1}y,\exp_{x}^{-1}z\rangle$
$\quad-2\theta\langle\exp_{z}^{-1}y,\exp_{z}^{-1}x\rangle$
$\leq\theta^{2}d^{2}(x,y)+d^{2}(x,z)+2\theta\|\exp_{x}^{-1}y-P_{x,z}\exp_{z}^{-1}y\|\|\exp_{x}^{-1}z\|$
$\quad-2\theta\langle\exp_{z}^{-1}y,\exp_{z}^{-1}x\rangle$
$\leq\theta^{2}d^{2}(x,y)+d^{2}(x,z)+2\theta d^{2}(x,z)-2\theta\langle\exp_{z}^{-1}y,\exp_{z}^{-1}x\rangle.$ (70)
From Lemma 40(ii), we have
$2\langle\exp_{z}^{-1}y,\exp_{z}^{-1}x\rangle\leq d^{2}(x,y)-d^{2}(y,z)-d^{2}(x,z).$ (71)
By combining (70) and (71), we get
$d^{2}(u,z)$ $\leq\theta^{2}d^{2}(x,y)+d^{2}(x,z)+2\theta d^{2}(x,z)+\theta d^{2}(x,y)-\theta d^{2}(y,z)-\theta d^{2}(x,z)$
$=(1+\theta)d^{2}(x,z)-\theta d^{2}(y,z)+\theta(1+\theta)d^{2}(x,y).$
Appendix B Proof of Proposition 15
For the proof of Proposition 15, we first recall [Leon2013, Proposition 4.1] for sectional curvature $k=0$.
Proposition 43 ([Leon2013]).
Let $\mathcal{M}$ be a Hadamard manifold with zero sectional curvature and let $x,y,x^{\prime},y^{\prime},a,b\in\mathcal{M}$ satisfy $x^{\prime}=\exp_{a}\bigl(-\exp_{a}^{-1}x\bigr)\quad\text{and}\quad y^{\prime}=\exp_{b}\bigl(-\exp_{b}^{-1}y\bigr).$ Suppose that $d(a,b)\leq d\bigl(\gamma(x,a;t),\gamma(y,b;t)\bigr)$ for all $t\in[0,1]$. Then $d(x^{\prime},y^{\prime})\leq d(x,y).$
Proof of Proposition 15. By the definition of the reflection, we have
$R_{\lambda f}(x)=\exp_{\operatorname*{Prox}_{\lambda f}(x)}\!\left(-\exp_{\operatorname*{Prox}_{\lambda f}(x)}^{-1}(x)\right),\quad R_{\lambda f}(y)=\exp_{\operatorname*{Prox}_{\lambda f}(y)}\!\left(-\exp_{\operatorname*{Prox}_{\lambda f}(y)}^{-1}(y)\right).$
Since $\operatorname*{Prox}_{\lambda f}$ is a firmly nonexpansive operator, it follows from (10) that
$d(\operatorname{Prox}_{\!\lambda f}(x),\operatorname{Prox}_{\!\lambda f}(y))\leq d(\gamma(x,\operatorname{Prox}_{\!\lambda f}(x);t),\gamma(y,\operatorname{Prox}_{\!\lambda f}(y);t))\text{ for all }t\in[0,1].$
Therefore, by Proposition 43, we obtain $d\bigl(R_{\lambda f}(x),R_{\lambda f}(y)\bigr)\leq d(x,y).$